### v2

# Modeling and computation of the effective elastic behavior of parallelogram origami metamaterials

Abstract
Origami metamaterials made of repeating unit cells of parallelogram panels joined at folds dramatically change their shape through a collective motion of their cells. Here we develop an effective elastic model and numerical method to study the large deformation response of these metamaterials under a broad class of loads. The model builds on an effective plate theory derived in our prior work [64]. The theory captures the overall shape change of all slightly stressed parallelogram origami deformations through nonlinear geometric compatibility constraints that couple the origami’s (cell averaged) effective deformation to an auxiliary angle field quantifying its cell-by-cell actuation. It also assigns to each such origami deformation a plate energy associated to these effective fields. Seeking a constitutive model that is faithful to the theory but also practical to simulate, we relax the geometric constraints via corresponding elastic energy penalties; we also simplify the plate energy density to embrace its essential character as a regularization to the geometric penalties. The resulting model for parallelogram origami is a generalized elastic continuum that is nonlinear in the effective deformation gradient and angle field and regularized by high-order gradients thereof. We provide a finite element formulation of this model using the $C^{0}$ interior penalty method to handle second gradients of deformation, and implement it using the open source computing platform Firedrake. We end by using the model and numerical method to study two canonical parallelogram origami patterns, in Miura and Eggbox origami, under a variety of loading conditions.
1 Introduction
Flexible mechanical metamaterials are cell-based patterns of stiff panels, connected at flexible hinges or folds, designed to achieve large overall shape-change at little overall stress. They are a promising class of materials for many applications, including for locomotion and grasping in soft robotics [30, 51, 65] and for the deployment of medical devices, space structures, and habitats [32, 41, 62, 71]. However, their mechanical behavior arises from a multiscale coupling spanning collective cell-wise interactions, large panel rotations, and localized distortion at the hinges or folds that is challenging to model and simulate. Here we study parallelogram origami — a large class of flexible mechanical metamaterials that exemplifies these challenges. Our goal is a modeling and computational framework that predicts the bulk macroscale response of these metamaterials under a broad range of loads.
Parallelogram origami is a family of mechanical metamaterials made of repeating unit cells of four parallelograms panels joined at folds. The most recognizable example is the Miura origami. Introduced as a means for packaging and deploying large space membranes decades ago [31], this pattern has since become an archetype for exploring functionality enabled by shape-morphing. It is the underlying motif in the inverse design of targeted shapes on morphing [14, 18, 53], can be used for self-folding at small and large scales [43, 59], and achieves auxetic compression and saddle-like bending modes under loads [55, 63]. Other well-studied examples of parallelogram origami include the Eggbox and Morph patterns [50, 56]. These patterns are distinct from the Miura in many ways. Both are (generically) non-Euclidean in that their vertex sector angles do not sum to $2\pi$. In addition, Eggbox has a positive Poisson’s ratio and is thus not an auxetic in contrast to the Miura. It also takes a cap-like shape when bent, rather than a saddle shape. The Morph, meanwhile, is distinguished by its versatility. It is either Miura-like in its mechanical behavior or Eggbox-like depending on the choice of mountain-valley assignment. Importantly, these patterns share one key mechanical property: Each can fold as a periodic mechanism [49] or floppy mode [35], i.e., a shape-changing continuous motion involving a periodic and rigid rearrangement of the panels about the flexible folds. It turns out that all parallelogram origami patterns are capable of folding as a periodic mechanism [39, 47, 64]. This property enables these patterns to dramatically change their shape under a wide range of loads, leading to a rich variety of soft modes of deformation (e.g., Fig. 1) whose stored elastic energy is far less than bulk, and inviting basic questions on how to best model their elastic behavior.
Figure 1: Examples of soft modes of deformation in paper models of Miura and Eggbox origami. (a-b) Pinching and bending Miura Origami. (c-d) Analogous examples for Eggbox origami.
By far the most popular modeling approach for these and other flexible mechanical metamaterials is bar-and-hinge elasticity [26, 33, 56] or related spring methods [15, 69]. These approaches replace the metamaterial with assemblies of elastic bars and hinges or introduce linear and torsional springs between the rigid elements, allowing the augmented material system to be analyzed using standard numerical methods in structural mechanics. Bar-and-hinge elasticity is particularly versatile and convenient for systems composed of a small number of building blocks, and the standard bearer in such settings [70]. However, it can become computationally expensive for large systems with many unit cells. It can also be challenging to fit the many spring stiffnesses in this method to yield accurate global behavior. Most importantly, the method does not provide a characterization of general soft modes of deformation, beyond simulating them. In particular, it does not explain why the Miura prefers a saddle shape when bent and the Eggbox prefers a cap.
An emerging alternative to bar-and-hinge elasticity is homogenization. The general idea is to replace the complex micro-motions of the metamaterial’s building blocks by an effective field theory that captures the material’s collective elastic interactions. The earliest success in this direction is perhaps the work of Alibert et al.[2] in their study of truss beams with pentographic substructure. Using the asymptotic variational method of $\Gamma$-convergence [7], they demonstrated that these structures coarse-grain to strain-gradient and higher-order elastic theories, assuming a linear elastic response. They later generalized these results to a myriad of other linear truss models of metamaterials, including origami and kirigami [1, 19], placing higher-order continuum theories on a rigorous foundation in this setting. A wide variety of researchers from engineering, mathematics and physics have added to this literature [4, 23, 28, 45, 54, 57, 61, 66], illustrating the richness and utility of coupling the metamaterial’s design to its effective elastic properties through homogenization. However, all of this research assumes up front that the metamaterial behaves linearly, and thus has only limited applicability to examples like parallelogram origami that dramatically change their shape.
We believe geometry should play the leading role when coarse-graining flexible metamaterials that possess mechanism deformations. For parallelogram origami, a basic physical heuristic is that their soft modes are locally mechanistic: at a lengthscale comparable to the size of a unit cell, each such mode looks like a mechanism. However, these mechanistic features can vary on the scale of many cells to produce a heterogenous actuation. The task then is to quantify exactly how the locally mechanistic response informs the overall shape. The first efforts in this direction are due to Schenk and Guest [55] and Wei et al.[63]. By studying the kinematics of fitting together slightly bent neighboring unit cells, both works uncovered a geometric link between the Miura’s in-­plane Poisson’s ratio and its normal curvatures, and used this link to explain why the Miura prefers a saddle shape when bent. Nassar et al.[46] built on these ideas by deriving the same Poisson’s ratio link for Eggbox origami. More perceptively, they viewed this link as encoding a global kinematic restriction on the effective surfaces possible in Eggbox origami, namely, by coupling it to the Gauss and Codazzi-Mainardi compatibility conditions from differential geometry [17]. Multiple lines of research have since demonstrated a fundamental coupling of Poisson’s ratio and normal curvatures in all parallelogram origami patterns [39, 47], in fact all periodic shells [44, 48], and connected this result to the origami’s effective shapes through differential geometry [12, 47]. Notably though, all of these results are a purely geometric attempt at capturing the soft modes observed in origami. They do not seek to systematically quantify the elastic energy of these modes through asymptotic analysis.
Our key theoretical contributions are to marry the nonlinear geometric constraints typical of flexible metamaterials with the systematic asymptotic analysis typical of linear homogenization. Our first work in this direction focused on the simpler 2D setting of planar kirigami [67]. There, we produced a coarse-graining rule that coupled the actuation of the kirigami’s panels and slits to its effective deformation through a metric constraint. We also justified this rule by showing that, for any solution to this metric constraint, there is a corresponding sequence of planar kirigami deformations converging to the given effective deformation, with elastic energy far less than bulk. We then turned our attention to parallelogram origami [64], which is more challenging due to the out-of-plane deformations that arise from panel bending. In brief, the coupling of effective deformation and cell-by-cell actuation is done through a metric constraint, similar to the kirigami setting, while the origami’s rigidity in bending constrains the curvature of the pattern to its Poisson’s ratio in line with prior results. These conditions can be expressed in a fully nonlinear way as two algebraic constraints coupling the first and second fundamental forms of the origami’s effective deformation to an angle field quantifying its cell-by-cell actuation (Eqs. (11) and (15) below). As our main technical achievement, we showed that every solution to these compatibility equations is the limit of a sequence of soft modes — origami deformations that involve large actuation of the folds, slight panel bending, and negligible stretching. Furthermore, we used these sequences to derive a plate theory for parallelogram origami patterns with an explicit coarse-grained quadratic energy depending on the second fundamental form of the effective deformation and the gradient of the actuation angle field. Section 2 provides an overview of this theory.
The purpose of the current work is to introduce a constitutive model for parallelogram origami that is faithful to the theory but also practical to simulate. A basic challenge is that it is hard to numerically implement the exact geometrical constraints that emerge when coarse-graining a flexible metamaterial. This is especially true under prototypical Dirichlet boundary conditions on effective deformation, as well as the tractions, moments and transverse loads typical of any engineering analysis of a plate. While mathematical efforts are under way by one of the co-authors to develop convergent numerical schemes that exactly solve such geometric constraints [36, 37, 38], the schemes are not yet capable of handling a wide variety of boundary conditions of practical interest. Here we embrace an alternative “relaxation and regularization” view that builds on the success of our work simulating the effective behavior of soft modes in planar kirigami [68], as well as other successful approaches to continuum modeling in the kirigami setting [11, 15, 40, 52]. Specifically, we relax the geometric constraints ((11) and (15)) via corresponding elastic energy penalties, and simplify the plate energy density from the theory to embrace its essential character as a regularization of the geometric penalties. The resulting model for parallelogram origami is a generalized elastic continuum reminiscent of, but distinct from, Eringen’s classical theories [21]. It is nonlinear in the effective deformation gradient and angle field and regularized by high-order gradients thereof. Importantly, it is also variationally well-posed (see Theorem 1), and thus ready-made for a numerical treatment based on finite elements. We provide a finite element formulation of the model using the $C^{0}$ interior penalty method, a well-established numerical technique for handling the higher-order gradients present in the model [20]. Implementation is then done via the open source computing platform Firedrake [29].
Our new continuum model and numerical method prove to be a versatile platform for solving elastic boundary value problems. This rings especially true in Section 5, where we showcase a compelling variety of simulations and analysis of the large deformation response of Miura and Eggbox origami under loads. One of the more interesting points exemplified by these examples is the strong coupling of the pattern’s Poisson’s ratio to its qualitative elastic behavior, beyond curvature. Our model produces equilibrium solutions where, depending on the boundary value problem, certain components of effective deformation or actuation are driven to approximate geometric partial differential equations (PDEs) that are either elliptic or hyperbolic, based solely on whether the pattern is auxetic or not. As elliptic and hyperbolic PDEs are dramatically different, so too is the behavior of Miura and Eggbox origami in their response to loads. Interestingly, the exact same connection between auxeticity and PDE type is also found in planar kirigami [67], suggesting perhaps that this characterization is a universal feature of the effective behavior of flexible mechanical metamaterials.
This paper is organized as follows. Section 2 describes the geometry of parallelogram origami and summarizes the effective theory derived in [64]. Section 3 introduces our generalized elastic continuum model, and its equilibrium equations and boundary conditions. Section 4 provides the finite element formulation and discusses its implementation. Section 5 deals with examples and Section 6 concludes the work.
2 Theoretical framework
We begin by recalling key ideas from [64], where we used asymptotic analysis to derive an effective continuum elastic plate theory for parallelogram origami. This theory forms the basis of the constitutive model and numerical method introduced later on to simulate the mechanical response of these patterns.
2.1 Designs and an effective description of origami mechanisms
Each parallelogram origami metamaterial is built from a repeating unit cell of four parallelograms panels joined at folds. As illustrated in Fig. 2(a) and (c), the design of such a pattern is fully parameterized by four design vectors $\mathbf{t}^{r}_{i}\in\mathbb{R}^{3}$, $i=1,2,3,4$, labeling the inner creases of a single unit cell in a counterclockwise fashion. The parallelogram panels are formed by pairs of opposite sides described by $\mathbf{t}^{r}_{i}$ and $\mathbf{t}^{r}_{i+1}$ (Fig. 2(a)). We assume throughout that these vectors correspond to a partially folded cell by enforcing the conditions
$\mathbf{t}^{r}_{i}\cdot(\mathbf{t}^{r}_{j}\times\mathbf{t}_{k}^{r})\neq 0\quad\text{ for all }ijk\in\{123,234,341,412\}.$ (1)
The cell is then tessellated along the Bravais lattice vectors
$\mathbf{u}_{0}:=\mathbf{t}_{1}^{r}-\mathbf{t}_{3}^{r},\quad\mathbf{v}_{0}:=\mathbf{t}_{2}^{r}-\mathbf{t}_{4}^{r}$ (2)
to produce an overall pattern (Fig. 2(c)). For future reference, we assume without loss of generality that $\mathbf{u}_{0}$ and $\mathbf{v}_{0}$ satisfy
$\mathbf{u}_{0}\cdot\mathbf{e}_{3}=\mathbf{v}_{0}\cdot\mathbf{e}_{3}=0,\quad\mathbf{e}_{3}\cdot(\mathbf{u}_{0}\times\mathbf{v}_{0})>0,$ (3)
and define their projection onto $\mathbb{R}^{2}$ as $\tilde{\mathbf{u}}_{0},\tilde{\mathbf{v}}_{0}$ such that $\mathbf{u}_{0}:=(\tilde{\mathbf{u}}_{0},0)$ and $\mathbf{v}_{0}:=(\tilde{\mathbf{v}}_{0},0)$. The design space of parallelogram origami is large. It includes many well-known examples like Miura origami [31], Eggbox origami [56] and the Morph patterns [50]. The theory we outline treats the design vectors as input, and thus encompasses all these patterns as special cases.
Figure 2: Parallelogram origami design and mechanism kinematics. (a) A parallelogram origami unit cell and the labeling of its creases. (b) Parameterization of its mechanism deformations. (c) The overall pattern is obtained by tessellating a unit cell. (d) The mechanism of the pattern repeats that of the unit cell.
A signature property of parallelogram origami is that each such pattern possesses a single degree-of-freedom (DOF) mechanism motion. Let us focus on the single cell in Fig. 2(a) with $\theta_{0}$ denoting the dihedral angle at the $\mathbf{t}^{r}_{1}$-crease. We fold this cell as origami by folding the crease from $\theta_{0}$ to $\theta_{0}+\theta$ while keeping the panels rigid. This actuation yields the deformed tangents in Fig. 2(b)
$\mathbf{t}^{r}_{i}\to\mathbf{t}_{i}(\theta),\quad\theta\in(\theta^{-},\theta^{+}),$ (4)
which satisfy $|\mathbf{t}_{i}(\theta)|=|\mathbf{t}_{i}^{r}|$ for each $i=1,\ldots,4$, and vary smoothly in $\theta$, thus describing a mechanism motion of the cell. [^1] Explicit parameterizations of this actuation are obtained by solving kinematic compatibility conditions for the panels to rigidly rotate about the folds of a four-fold vertex. Such parameterizations are given in [25, 58] for rigidly and flat foldable origami and in [27] for the general Euclidean and non-Euclidean origami. The interval $(\theta^{-},\theta^{+})$ depends only on $\mathbf{t}_{1}^{r},\ldots,\mathbf{t}_{4}^{r}$, and is characterized as the largest open interval containing zero that satisfies
$\big{[}\mathbf{t}^{r}_{i}\cdot(\mathbf{t}^{r}_{j}\times\mathbf{t}_{k}^{r})\big{]}\big{[}\mathbf{t}_{i}(\theta)\cdot(\mathbf{t}_{j}(\theta)\times\mathbf{t}_{k}(\theta))\big{]}>0\quad\text{ for all }ijk\in\{123,234,341,412\}.$ (5)
Such an interval exists for any parallelogram origami design satisfying (1). [^2] A proof of this fact is given in [64]; it uses an argument based on the implicit function theorem. The overall pattern in Fig. 2(c), in turn, has a mechanism motion that repeats the motion of the cell (Fig. 2(d)). It is obtained by tessellating a deformed cell along the Bravais lattice vectors
$\mathbf{u}(\theta):=\mathbf{t}_{1}(\theta)-\mathbf{t}_{3}(\theta),\quad\mathbf{v}(\theta):=\mathbf{t}_{2}(\theta)-\mathbf{t}_{4}(\theta),\quad\theta\in(\theta^{-},\theta^{+}).$ (6)
As in the reference state, we assume without loss of generality that
$\mathbf{u}(\theta)\cdot\mathbf{e}_{3}=\mathbf{v}(\theta)\cdot\mathbf{e}_{3}=0,\quad\mathbf{e}_{3}\cdot(\mathbf{u}(\theta)\times\mathbf{v}(\theta))>0$ (7)
to express the basic fact that the mechanism motion is effectively planar. Furthermore, we define the shape tensor as the unique linear transformation $\mathbf{A}(\theta)\in\mathbb{R}^{3\times 2}$ such that
$\mathbf{A}(\theta)\tilde{\mathbf{u}}_{0}=\mathbf{u}(\theta),\quad\mathbf{A}(\theta)\tilde{\mathbf{v}}_{0}=\mathbf{v}(\theta).$ (8)
Note that $\mathbf{e}_{3}\cdot\mathbf{A}(\theta)=\mathbf{0}$ per (7).
2.2 Coarse-graining rules and an effective description of origami soft modes
Figure 3: Illustration of the locally mechanistic coarse-graining rule. (a) A finely patterned parallelogram origami, with cells of size $\sim\ell$, covers a unit square domain. (b) A soft mode of this pattern corresponds to a spatially varying local mechanism.
Beyond the pure mechanisms, parallelogram origami can bend, twist and actuate in complex nonuniform ways at very little elastic energy by localizing most of the deformation at the folds. We call such deformations soft modes. As anyone who has folded a Miura origami knows, soft modes are the generic response of these patterns to loads. Here, we describe a theory that captures the effective behavior of all possible soft modes in parallelogram origami. As explained below, it is built on two coarse-graining rules that link the design and kinematics of the origami cells explicitly to its effective cell-averaged deformation.
The first coarse-graining rule originates from our earlier work on planar kirigami [67, 68]. Soft modes are locally mechanistic. At the scale of each unit cell they look like a mechanism. However, their features vary slowly from cell to cell. This rule reveals a fundamental relationship between the first fundamental form of the effective deformation $\mathbf{y}(\mathbf{x})$ of a soft mode
$\mathbf{I}(\mathbf{y}):=(\nabla\mathbf{y})^{T}\nabla\mathbf{y}$ (9)
and the cell-by-cell actuation $\theta=\theta(\mathbf{x})$, which we now think of as a spatially varying field. Fig. 3 illustrates these ideas by way of a concrete example. A reference parallelogram origami pattern is overlaid onto a square domain $\Omega=(0,1)^{2}$ with unit cells now of characteristic length $\sim\ell$ in Fig. 3(a) and subject to a soft mode of deformation in Fig. 3(b). Notice that, while the origami surface oscillates due to the underlying microstructure of the unit cell, corresponding vertices in the reference pattern (the ones related by a translation $m\ell\mathbf{u}_{0}+n\ell\mathbf{v}_{0}$, $m,n\in\mathbb{Z}$) evolve smoothly from cell to cell. Since the soft mode also exhibits the same corrugated features as the mechanism, this evolution is quantified by an underlying effective deformation $\mathbf{y}\colon\Omega\rightarrow\mathbb{R}^{3}$, rotation field $\mathbf{R}\colon\Omega\rightarrow SO(3)$, and actuation field $\theta\colon\Omega\rightarrow(\theta^{-},\theta^{+})$ that satisfy (see Fig. 3)
$\mathbf{y}(\mathbf{x}+\ell\tilde{\mathbf{u}}_{0})-\mathbf{y}(\mathbf{x})\approx\ell\partial_{\mathbf{u}_{0}}\mathbf{y}(\mathbf{x})\approx\ell\mathbf{R}(\mathbf{x})\mathbf{u}(\theta(\mathbf{x})),$ (10)
$\mathbf{y}(\mathbf{x}+\ell\tilde{\mathbf{v}}_{0})-\mathbf{y}(\mathbf{x})\approx\ell\partial_{\mathbf{v}_{0}}\mathbf{y}(\mathbf{x})\approx\ell\mathbf{R}(\mathbf{x})\mathbf{v}(\theta(\mathbf{x})).$
It immediately follows that the fields $\theta$ and $\mathbf{I}(\mathbf{y})$ are coupled through the shape tensor in (8):
$\mathbf{I}(\mathbf{y})=\mathbf{A}^{T}(\theta)\mathbf{A}(\theta).$ (11)
While most of the deformation is localized at the folds in a soft mode, the panels must bend slightly to accommodate a non-uniform actuation from cell to cell. Our second coarse-graining rule is to constrain the effective theory by how neighboring unit cells with slightly bent panels fit together. This rule reveals a fundamental relationship between the actuation $\theta(\mathbf{x})$ and the second fundamental form of the effective deformation
$\mathbf{II}(\mathbf{y}):=\begin{pmatrix}\partial_{1}\partial_{1}\mathbf{y}\cdot\mathbf{n}(\mathbf{y})&\partial_{1}\partial_{2}\mathbf{y}\cdot\mathbf{n}(\mathbf{y})\\ \partial_{1}\partial_{2}\mathbf{y}\cdot\mathbf{n}(\mathbf{y})&\partial_{2}\partial_{2}\mathbf{y}\cdot\mathbf{n}(\mathbf{y})\end{pmatrix}\quad\text{ for }\quad\mathbf{n}(\mathbf{y}):=\frac{\partial_{1}\mathbf{y}\times\partial_{2}\mathbf{y}}{|\partial_{1}\mathbf{y}\times\partial_{2}\mathbf{y}|}.$ (12)
Fig. 4 illustrates the idea. Zoom into the local neighborhood of cells in a soft mode and identify the rotation of the base cell by $\mathbf{R}(\mathbf{x})$, as indicated, to be consistent with (LABEL:eq:keyCoarseGrain1) and Fig. 3. Given the slightly bent nature of all the panels in this description, the cells rotate relative to each other by a small amount. We quantify this kinematics at leading order by noticing that the rotation of the cell to the right and above the base cell satisfy
$\mathbf{R}(\mathbf{x}+\ell\tilde{\mathbf{u}}_{0})\approx\mathbf{R}(\mathbf{x})\big{(}\mathbf{I}+\ell(\bm{\omega}_{\mathbf{u}_{0}}(\mathbf{x})\times)\big{)},$ (13)
$\mathbf{R}(\mathbf{x}+\ell\tilde{\mathbf{v}}_{0})\approx\mathbf{R}(\mathbf{x})\big{(}\mathbf{I}+\ell(\bm{\omega}_{\mathbf{v}_{0}}(\mathbf{x})\times)\big{)},$
respectively, for some vectors fields $\bm{\omega}_{\mathbf{u}_{0}},\bm{\omega}_{\mathbf{v}_{0}}\colon\Omega\rightarrow\mathbb{R}^{3}$ that measure bending in the pattern. [^3] Note, $(\mathbf{v}\times)$ denotes a tensor on $\mathbb{R}^{3\times 3}$ that satisfies $(\mathbf{v}\times)\mathbf{w}=\mathbf{v}\times\mathbf{w}$ for all $\mathbf{w}\in\mathbb{R}^{3}$ for any $\mathbf{v}\in\mathbb{R}^{3}$; in other words, $(\mathbf{v}\times)$ is a generic parameterization of a skew tensor. Since the cells must fit together at their boundaries, the fields $\bm{\omega}_{\mathbf{u}_{0}}$, $\bm{\omega}_{\mathbf{v}_{0}}$ and $\theta$ are not independent of each other. It turns out that they satisfy (see [64])
$\bm{\omega}_{\mathbf{u}_{0}}\cdot\mathbf{v}^{\prime}(\theta)=\bm{\omega}_{\mathbf{v}_{0}}\cdot\mathbf{u}^{\prime}(\theta).$ (14)
As a final bit of manipulation, this constraint can be rewritten using the implied identity $\partial_{\mathbf{u}_{0}}\big{[}\mathbf{R}\mathbf{v}(\theta)\big{]}=\partial_{\mathbf{v}_{0}}\big{[}\mathbf{R}\mathbf{u}(\theta)\big{]}$ from (LABEL:eq:keyCoarseGrain1). Doing so eliminates the bending measures $\bm{\omega}_{\mathbf{u}_{0}}$, $\bm{\omega}_{\mathbf{v}_{0}}$ in favor of the second fundamental form $\mathbf{II}(\mathbf{y})$, giving an equivalent constraint in terms of the fields $\mathbf{y}$ and $\theta$:
$\big{[}\mathbf{v}(\theta)\cdot\mathbf{v}^{\prime}(\theta)\big{]}\big{[}\tilde{\mathbf{u}}_{0}\cdot\mathbf{II}(\mathbf{y})\tilde{\mathbf{u}}_{0}\big{]}+\big{[}\mathbf{u}(\theta)\cdot\mathbf{u}^{\prime}(\theta)\big{]}\big{[}\tilde{\mathbf{v}}_{0}\cdot\mathbf{II}(\mathbf{y})\tilde{\mathbf{v}}_{0}\big{]}=0.$ (15)
Figure 4: Local fitting problem for slightly bent unit cells. Neighboring cells rotate relative to each other due to panel bending. These rotations are coupled through the compatibility conditions at the cell boundaries.
2.3 Asymptotic analysis and an effective plate theory
The bulk of our effort in [64] goes into justifying the constraints in (11) and (15). We showed that every sufficiently smooth solution to these effective compatibility equations is the limit of a sequence of soft modes — origami deformations with negligible panel strain that approximate the effective deformation. The precise result is as follows: Let $\Omega\subset\mathbb{R}^{2}$ be a simply connected domain of characteristic length $\sim 1$, and let $\{\mathbf{x}_{i,j}^{\ell}\}\subset\mathbb{R}^{3}$ label the vertices of a reference parallelogram origami domain overlaid onto $\Omega$ whose cells are of characteristic length $\ell\ll 1$ (see, e.g., Fig. 3). Then, for any sufficiently smooth $\theta\colon\Omega\rightarrow(\theta^{-},\theta^{+})$ and $\mathbf{y}\colon\Omega\rightarrow\mathbb{R}^{3}$ that solve (11) and (15), there is a corresponding deformation of the origami vertices $\{\mathbf{x}_{i,j}^{\ell}\}\rightarrow\{\mathbf{y}_{i,j}^{\ell}\}$ such that
$\frac{|\mathbf{y}_{i,j}^{\ell}-\mathbf{y}_{i+1,j}^{\ell}|-|\mathbf{x}_{i,j}^{\ell}-\mathbf{x}_{i+1,j}^{\ell}|}{|\mathbf{x}_{i,j}^{\ell}-\mathbf{x}_{i+1,j}^{\ell}|}+\frac{|\mathbf{y}_{i,j}^{\ell}-\mathbf{y}_{i,j+1}^{\ell}|-|\mathbf{x}_{i,j+1}^{\ell}-\mathbf{x}_{i,j+1}^{\ell}|}{|\mathbf{x}_{i,j}^{\ell}-\mathbf{x}_{i,j+1}^{\ell}|}$ (16)
$\qquad\qquad+\frac{|\mathbf{y}_{i,j}^{\ell}-\mathbf{y}_{i+1,j}^{\ell}|-|\mathbf{x}_{i,j}^{\ell}-\mathbf{x}_{i+1,j}^{\ell}|}{|\mathbf{x}_{i,j}^{\ell}-\mathbf{x}_{i+1,j}^{\ell}|}+\frac{|\mathbf{y}_{i,j}^{\ell}-\mathbf{y}_{i,j+1}^{\ell}|-|\mathbf{x}_{i,j+1}^{\ell}-\mathbf{x}_{i,j+1}^{\ell}|}{|\mathbf{x}_{i,j}^{\ell}-\mathbf{x}_{i,j+1}^{\ell}|}=O(\ell^{2}),$
$|\mathbf{y}(\tilde{\mathbf{x}}_{i,j}^{\ell})-\mathbf{y}_{i,j}^{\ell}|=O(\ell)$
for all $(i,j)$, where $\tilde{\mathbf{x}}_{i,j}^{\ell}$ is the projection of $\mathbf{x}_{i,j}^{\ell}$ onto $\Omega$. The first statement in (LABEL:eq:mainResult) says that the strain of every panel is at most $\sim\ell^{2}$. The second says that the origami deformation approximates the effective deformation.
In fact, we can say more. Consider a protypical bar-and-hinge energy [26, 33] composed of stretching, bending and folding terms, denoted $E_{\text{tot}}^{\ell}(\{\mathbf{y}_{i,j}^{\ell}\})=E^{\ell}_{\text{str}}(\{\mathbf{y}_{i,j}^{\ell}\})+E^{\ell}_{\text{bend}}(\{\mathbf{y}_{i,j}^{\ell}\})+E^{\ell}_{\text{fold}}(\{\mathbf{y}_{i,j}^{\ell}\})$, and scale the moduli of these three terms by powers of $\ell$ such that “stretching stiffness” $\gg$ “bending stiffness” $\gg$ “folding stiffness”. With the help of the constructions above, we obtain the limit of this energy
$\lim_{\ell\rightarrow 0}E^{\ell}_{\text{tot}}(\{\mathbf{y}_{i,j}^{\ell}\})=\begin{cases}\int_{\Omega}\begin{pmatrix}\mathbf{II}(\mathbf{y})\\ (\nabla\theta)^{T}\end{pmatrix}\colon\mathbb{K}(\theta)\colon\begin{pmatrix}\mathbf{II}(\mathbf{y})\\ (\nabla\theta)^{T}\end{pmatrix}\mathop{}\!\mathrm{d}x&\text{ if }(\theta,\mathbf{y})\in A_{\text{eff}}\\ +\infty&\text{ otherwise,}\end{cases}$ (17)
where $\mathbb{K}(\theta)\in\mathbb{R}^{3\times 2\times 3\times 2}$ is an explicit fourth order tensor with major symmetry $[\mathbb{K}(\theta)]_{i\alpha j\beta}=[\mathbb{K}(\theta)]_{j\beta i\alpha}$ that depends kinematically only on the actuation field $\theta$ (see [64] for the full definition). The set $A_{\text{eff}}$ denotes the class of admissible effective deformations and actuation fields describing a soft mode
[…]
To derive this plate theory, we focused on smooth effective fields within $A_{\text{eff}}$. For any such fields, we showed that the total energy of the corresponding origami construction $\{\mathbf{y}_{i,j}^{\ell}\}$ is dominated by panel bending $E_{\text{tot}}^{\ell}(\{\mathbf{y}_{i,j}^{\ell}\})\approx E_{\text{bend}}^{\ell}(\{\mathbf{y}_{i,j}^{\ell}\})\sim 1$, yielding the finite limit in the first part of the statement in (17). However, when the effective fields depart this set, the panel strain of the constructions is $O(\ell)$ or larger, as opposed to the $O(\ell^{2})$ (see (LABEL:eq:mainResult)). The total energy is then dominated by panel stretching $E_{\text{tot}}^{\ell}(\{\mathbf{y}_{i,j}^{\ell}\})\approx E_{\text{str}}^{\ell}(\{\mathbf{y}_{i,j}^{\ell}\})\gg 1$, which is unbounded in the limit $\ell\rightarrow 0$ as indicated by the “+$\infty$+ ∞” in (17). The main point is that soft modes in parallelogram origami are characterized by effective fields that belong to $A_{\text{eff}}$ and possess a quadratic energy density in $\mathbf{II}(\mathbf{y})$ and $\nabla\theta$.
3 Modeling framework
Building on the theory, we now develop a constitutive model well-adapted to simulate the effective behavior of soft modes in parallelogram origami. A basic challenge is that it is hard to numerically implement exact geometrical constraints, like the ones in $A_{\text{eff}}$. This is especially true under prototypical displacement and traction boundary conditions on the effective deformation $\mathbf{y}$. We seek instead a practical continuum constitutive model that is faithful to the theory but also efficient to simulate and predictive.
3.1 Constitutive model
We take as our effective constitutive model the sum of bulk and regularizing terms
$E_{\text{int}}(\mathbf{y},\theta):=E_{\text{bulk}}(\mathbf{y},\theta)+E_{\text{reg}}(\mathbf{y},\theta),$ (19)
measuring the stored elastic energy due to the effective deformation and actuation of a parallelogram origami pattern on an underlying reference domain $\Omega\subset\mathbb{R}^{2}$.
We choose the bulk energy in this formulation to relax the geometric constraints in $A_{\text{eff}}$ via
$E_{\text{bulk}}(\mathbf{y},\theta)$ $:=\int_{\Omega}\Big{\{}W_{1}(\theta,\mathbf{I}(\mathbf{y}))+W_{2}(\theta,\mathbf{II}(\mathbf{y}))\Big{\}}\mathop{}\!\mathrm{d}x.$ (20)
In particular, the two energy densities $W_{1,2}\colon(\theta^{-},\theta^{+})\times\mathbb{R}_{\text{sym}}^{2\times 2}\rightarrow\mathbb{R}$ are penalizations of the constraints in (11) and (15) defined by
$W_{1}(\theta,\mathbf{G}):=\frac{c_{1}}{\sqrt{\det\mathbf{G}}}\Big{|}\mathbf{G}-\mathbf{A}^{T}(\theta)\mathbf{A}(\theta)\Big{|}^{2},$ (21)
$W_{2}(\theta,\mathbf{K}):=\frac{c_{2}L_{\Omega}^{2}}{|\tilde{\mathbf{u}}_{0}|^{4}|\tilde{\mathbf{v}}_{0}|^{4}}\Big{(}\big{[}\mathbf{v}(\theta)\cdot\mathbf{v}^{\prime}(\theta)\big{]}\big{[}\tilde{\mathbf{u}}_{0}\cdot\mathbf{K}\tilde{\mathbf{u}}_{0}\big{]}|+\big{[}\mathbf{u}(\theta)\cdot\mathbf{u}^{\prime}(\theta)\big{]}\big{[}\tilde{\mathbf{v}}_{0}\cdot\mathbf{K}\tilde{\mathbf{v}}_{0}\big{]}\Big{)}^{2}.$
The division by $\sqrt{\det\mathbf{I}(\mathbf{y})}$ in the $W_{1}(\theta,\mathbf{I}(\mathbf{y}))$ seeks to prevent interpenetration of matter in the model, while the division by $|\tilde{\mathbf{u}}_{0}|^{4}|\tilde{\mathbf{v}}_{0}|^{4}$ in $W_{2}(\theta,\mathbf{II}(\mathbf{y}))$ scales this energy density so that it is invariant under conformal transformations of the unit cell, i.e., given by $(\tilde{\mathbf{u}}_{0},\tilde{\mathbf{v}}_{0})\rightarrow\lambda(\tilde{\mathbf{u}}_{0},\tilde{\mathbf{v}}_{0})$ and $(\mathbf{u}(\theta),\mathbf{v}(\theta))\rightarrow\lambda(\mathbf{u}(\theta),\mathbf{v}(\theta))$ for all $\lambda>0$. Finally, $L_{\Omega}>0$ denotes a characteristic lengthscale of the domain $\Omega$. It is introduced in $W_{2}$ so that $c_{2}$ has dimensions of energy per unit area, just like $c_{1}$.
Having captured the key nonlinearities in the bulk term, we invoke the adage that “simplicity is best” for the regularizing term by defining
$E_{\text{reg}}(\mathbf{y},\theta):=\int_{\Omega}\Big{\{}d_{1}L_{\Omega}^{2}|\nabla\nabla\mathbf{y}|^{2}+d_{2}L_{\Omega}^{2}|\nabla\theta|^{2}+d_{3}\theta^{2}\Big{\}}\mathop{}\!\mathrm{d}x.$ (22)
The first and second energy densities in (22) are simple expressions that serve to both approximate the plate energy density in (17) and regularize the bulk energy in (20). They involve higher order gradients and thus, like $W_{2}$, are scaled by $L_{\Omega}^{2}$ to set the dimensions of $d_{1},d_{2}$ as energy per unit area. The last term is the simplest effective energy density that accounts for the higher order but ubiquitous elastic energy to fold origami.
Note that, while we eventually anticipate fitting $c_{1},c_{2},d_{1},d_{2},d_{3}>0$ to experiments, our asymptotic analysis [64] suggests that these moduli should obey the relations
$c_{1}\gg c_{2}\gg d_{1}\sim d_{2}$ (23)
in the typical case. We also assumed $d_{3}\ll c_{1},\ldots,d_{2}$ in [64] to simplify the analysis. However, the appropriate choice of this folding modulus really depends in a delicate way on the stiffness ratio between the panels and folds. In Section 5, for instance, we choose $d_{3}$ to be larger than $d_{1,2}$ in most examples, as this choice produces simulations that conform well to the paper models of the origami in Fig. 1; $d_{3}$ should always be $\ll c_{1},c_{2}$.
As a final point on the constitutive model, we take the the ambient space on which to define and study this energy to be
$V^{\varepsilon}:=\big{\{}(\mathbf{y},\theta)\in H^{2}(\Omega,\mathbb{R}^{3})\times H^{1}(\Omega,\mathbb{R})\colon|\partial_{1}\mathbf{y}\times\partial_{2}\mathbf{y}|\geq\varepsilon\text{ and }\theta\in[\theta^{-}+\varepsilon,\theta^{+}-\varepsilon]\text{ a.e.\;in }\Omega\big{\}}$ (24)
for a small positive parameter $\varepsilon>0$. Since $\nabla\nabla\mathbf{y}$ and $\nabla\theta$ must be square-integrable for $E_{\text{int}}(\mathbf{y},\theta)$ to be well-defined and bounded, restricting our attention to deformations and actuation fields in $H^{2}\times H^{1}$ is well-justified. The $\varepsilon$-inequalities are not strictly needed to obtain a bounded energy, but are physically motivated and help to simplify some technical issues in proving the well-posedness of the model. For instance, the function $\mathbf{u}(\theta)$ and $\mathbf{v}(\theta)$ are well-defined and smooth on the interval $(\theta^{-},\theta^{+})$ but can become poorly behaved as $\theta\rightarrow\theta^{\pm}$. In fact, there are additional physics at play in these limits, corresponding to a change in mountain-valley assignment, that we do not account for in our modeling. The restriction $\theta\in[\theta^{-}+\varepsilon,\theta^{+}-\varepsilon]$ rules out this behavior. The other inequality $|\partial_{1}\mathbf{y}\times\partial_{2}\mathbf{y}|\geq\varepsilon$ ensures that the surface normal $\mathbf{n}(\mathbf{y})$ and $1/\sqrt{\det\mathbf{I}(\mathbf{y})}$ are always well-defined and, in particular, that the admissible class of effective surfaces are bounded away from ones that contain singularities. While both restrictions are crucial to our proof of the well-posedness of the model, they have no impact on how we numerically implement the model in practice.
3.2 Boundary conditions and applied forces
We investigate the effective behavior of parallelogram origami under Dirichlet boundary conditions, corresponding to applied deformation and slope, as well as applied forces, generalized moments, and distributed loads. Let $\Gamma_{\text{d}},\Gamma_{\text{n}}\subset\partial\Omega$ denote relatively open subsets of the boundary of $\Omega$. We prescribe the deformation and slope, respectively, via
$\mathbf{y}=\overline{\mathbf{y}}\text{ on }\Gamma_{\text{d}}\quad\text{ and }\quad(\nabla\mathbf{y})\mathbf{n}=\overline{\mathbf{s}}\text{ on }\Gamma_{\text{n}}$ (25)
for some $\overline{\mathbf{y}}\colon\Gamma_{\text{d}}\rightarrow\mathbb{R}^{3}$ and $\overline{\mathbf{s}}\colon\Gamma_{\text{n}}\rightarrow\mathbb{R}^{3}$.
To incorporate applied forces, we invoke the principle of minimum potential energy. The total potential energy is the sum of the stored elastic energy in (19) and the contribution associated to applied forces, written here as
$E(\mathbf{y},\theta):=E_{\text{int}}(\mathbf{y},\theta)+E_{\text{ext}}(\mathbf{y}).$ (26)
Let $\Gamma_{\text{t}}:=\partial\Omega\setminus\Gamma_{\text{d}}$ and $\Gamma_{\text{m}}:=\partial\Omega\setminus\Gamma_{\text{n}}$ denote the boundary sets on which we have not applied the deformation or slope. We allow for applied forces on these sets via the external energy potential
$E_{\text{ext}}(\mathbf{y}):=-\int_{\Omega}\overline{\mathbf{b}}\cdot\mathbf{y}\mathop{}\!\mathrm{d}x-\int_{\Gamma_{\text{t}}}\overline{\mathbf{t}}\cdot\mathbf{y}\mathop{}\!\mathrm{d}s-\int_{\Gamma_{\text{m}}}\overline{\mathbf{m}}\cdot(\nabla\mathbf{y})\mathbf{n}\mathop{}\!\mathrm{d}s,$ (27)
where $\overline{\mathbf{b}}\colon\Omega\rightarrow\mathbb{R}^{3}$ denotes the prescribed referential distributed loads, $\overline{\mathbf{t}}\colon\Gamma_{\text{t}}\rightarrow\mathbb{R}^{3}$ the prescribed referential traction, and $\overline{\mathbf{m}}\colon\Gamma_{\text{m}}\rightarrow\mathbb{R}^{3\times 2}$ the prescribed referential generalized moment.
Note that the boundary conditions correspond only to $\mathbf{y}$, $\nabla\mathbf{y}$ and their pair-conjugate generalized forces. In principle, one could also include a Dirichlet boundary conditions on $\theta$ and its pair-conjugate torque. We avoid this inclusion for simplicity and because such boundary conditions seem hard to realize experimentally in parallelogram origami. [^4] These boundary conditions are, however, relevant to planar kirigami metamaterials. A nice example comes from [15], where they used “slit actuation” boundary conditions to enable a domain wall in experiments on these systems.
3.3 Energy minimization and an existence theorem
Having defined the model, boundary conditions, and forces, we now establish the existence of minimizers of the energy. Let
$E^{\star}:=\inf\big{\{}E(\mathbf{y},\theta)\colon(\mathbf{y},\theta)\in V^{\varepsilon}_{\Gamma}\big{\}}$ (28)
for $V^{\varepsilon}_{\Gamma}$ the subset of effective deformations and actuation fields in $V^{\varepsilon}$ that enforces the desired Dirichlet boundary conditions
$V^{\varepsilon}_{\Gamma}:=\big{\{}(\mathbf{y},\theta)\in V^{\varepsilon}\colon\mathbf{y}=\overline{\mathbf{y}}\text{ a.e.\;on }\Gamma_{\text{d}}\text{ and }(\nabla\mathbf{y})\mathbf{n}=\overline{\mathbf{s}}\text{ a.e.\;on }\Gamma_{\text{n}}\big{\}}.$ (29)
The equalities at the boundaries $\Gamma_{\text{d}}$ and $\Gamma_{\text{n}}$ are to be understood in the sense of traces. [^5] In particular, we assume that $\overline{\mathbf{y}}\in H^{\frac{1}{2}}(\Gamma_{\text{d}},\mathbb{R}^{3})$ and $\bar{\mathbf{s}}\in H^{\frac{1}{2}}(\Gamma_{\text{n}},\mathbb{R}^{3})$, as these are the appropriate trace Sobolev spaces for such boundary conditions (see Evans [24], Chapter 5, and Brezis [8], Chapter 9, for more details). We have the following result.
Theorem 1.
Assume that $\Gamma_{\emph{n}}\subset\Gamma_{\emph{d}}$ are measurable, $\Gamma_{\emph{d}}$ has nonzero measure, and that $V^{\varepsilon}_{\Gamma}$ is non-empty. Let $\overline{\mathbf{b}}\in L^{2}(\Omega,\mathbb{R}^{3})$, $\overline{\mathbf{t}}\in L^{2}(\Gamma_{\emph{t}},\mathbb{R}^{3})$, $\overline{\mathbf{m}}\in L^{2}(\Gamma_{\emph{m}},\mathbb{R}^{3\times 3})$. Then, there is a $(\mathbf{y},\theta)\in V^{\varepsilon}_{\Gamma}$ such that $E^{\star}=E(\mathbf{y},\theta)$.
A proof of this theorem is provided in Appendix A. The key point is that the energy $E(\mathbf{y},\theta)$, $(\mathbf{y},\theta)\in V_{\Gamma}^{\varepsilon}$, is weakly lower semicontintinuous, a result that follows from a general classification of lower semicontinuous functionals in Dacorogna [13], Theorem 3.23. There is some subtlety to this point. The hypotheses needed to apply the Dacorogna result are not directly compatible with the $\varepsilon$-inequalities of $V_{\Gamma}^{\varepsilon}$. However, we are able to modify the energy in order to apply the theorem to our setting and then argue that the modification is, in fact, benign. The rest of the proof is more-or-less standard fare in the calculus of variation. We provide the details in the appendix for the benefit of the mechanics reader who is less familiar with the techniques.
3.4 Weak formulation and equilibrium equations
We now derive the equilibrium equations for our model by taking the first variation of the energy. For this purpose, we dispense with the $\varepsilon$-dependence of the space of admissible fields in (29), as it is not needed to develop this result, and instead consider the space
$V_{\Gamma}:=\big{\{}(\mathbf{y},\theta)\in H^{2}(\Omega,\mathbb{R}^{3})\times H^{1}(\Omega,(\theta^{-},\theta^{+}))\colon\mathbf{y}=\overline{\mathbf{y}}\text{ a.e.\;on }\Gamma_{\text{d}}\text{ and }(\nabla\mathbf{y})\mathbf{n}=\overline{\mathbf{s}}\text{ a.e.\;on }\Gamma_{\text{n}}\big{\}}.$ (30)
We also denote by $V_{0}$ the space of functions in (30) such that $\overline{\mathbf{y}}=\mathbf{0}$ and $\overline{\mathbf{s}}=\mathbf{0}$.
Let $(\mathbf{y},\theta)\in V_{\Gamma}$ be a local minimizer to the energy $E$. The first variation of this energy satisfies
00 0 $=\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}E(\mathbf{y}+t\mathbf{w},\theta+t\eta)\big{|}_{t=0}$ (31)
$=\int_{\Omega}\Big{\{}\mathbf{P}(\mathbf{y},\theta)\colon\nabla\mathbf{w}+\big{\langle}\mathcal{H}(\mathbf{y},\theta),\nabla\nabla\mathbf{w}\big{\rangle}+q(\mathbf{y},\theta)\eta+\mathbf{j}(\theta)\cdot\nabla\eta\Big{\}}\mathop{}\!\mathrm{d}x$
$\qquad-\int_{\Omega}\overline{\mathbf{b}}\cdot\mathbf{w}\mathop{}\!\mathrm{d}x-\int_{\Gamma_{\text{t}}}\overline{\mathbf{t}}\cdot\mathbf{w}\mathop{}\!\mathrm{d}s-\int_{\Gamma_{\text{m}}}\overline{\mathbf{m}}\cdot(\nabla\mathbf{w})\mathbf{n}\mathop{}\!\mathrm{d}s$
[…]
$\mathbf{P}(\mathbf{y},\theta):=\mathbf{P}_{1}(\mathbf{y},\theta)+\mathbf{P}_{2}(\mathbf{y},\theta)$ (32)
for second-order tensor fields $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ on $\mathbb{R}^{3\times 2}$ obtained by differentiating the energy densities $W_{1}$ and $W_{2}$ with respect to the deformation gradient $\nabla\mathbf{y}$. This differentiation furnishes
$\mathbf{P}_{1}(\mathbf{y},\theta):=\frac{4c_{1}}{\sqrt{\det\mathbf{I}(\mathbf{y})}}(\nabla\mathbf{y})\Big{[}\big{(}\mathbf{I}(\mathbf{y})-\mathbf{A}^{T}(\theta)\mathbf{A}(\theta)\big{)}-\frac{1}{4}|\mathbf{I}(\mathbf{y})-\mathbf{A}^{T}(\theta)\mathbf{A}(\theta)|^{2}\big{(}\mathbf{I}(\mathbf{y})\big{)}^{-1}\Big{]}$ (33)
$\mathbf{P}_{2}(\mathbf{y},\theta):=\frac{2c_{2}L_{\Omega}^{2}k(\mathbf{y},\theta)}{|\tilde{\mathbf{u}}_{0}|^{4}|\tilde{\mathbf{v}}_{0}|^{4}|\partial_{1}\mathbf{y}\times\partial_{2}\mathbf{y}|}\Big{[}(\partial_{2}\mathbf{y}\times\mathbf{N}(\mathbf{y})L(\theta)\mathbf{y})\otimes\mathbf{e}_{1}-(\partial_{1}\mathbf{y}\times\mathbf{N}(\mathbf{y})L(\theta)\mathbf{y})\otimes\mathbf{e}_{2}\Big{]},$
after some algebraic manipulation. The latter formula includes $k(\mathbf{y},\theta):=\big{[}\mathbf{v}(\theta)\cdot\mathbf{v}^{\prime}(\theta)\big{]}\big{[}\tilde{\mathbf{u}}_{0}\cdot\mathbf{II}(\mathbf{y})\tilde{\mathbf{u}}_{0}\big{]}+\big{[}\mathbf{u}(\theta)\cdot\mathbf{u}^{\prime}(\theta)\big{]}\big{[}\tilde{\mathbf{v}}_{0}\cdot\mathbf{II}(\mathbf{y})\tilde{\mathbf{v}}_{0}\big{]}$, which relaxes the left side of the kinematic constraint in (15). It also includes the projection tensor $\mathbf{N}(\mathbf{y}):=\mathbf{I}-\mathbf{n}(\mathbf{y})\otimes\mathbf{n}(\mathbf{y})$, and the linear operator $L(\theta):=[(\mathbf{v}(\theta)\cdot\mathbf{v}^{\prime}(\theta))\partial_{\mathbf{u}_{0}}\partial_{\mathbf{u}_{0}}+(\mathbf{u}(\theta)\cdot\mathbf{u}^{\prime}(\theta))\partial_{\mathbf{v}_{0}}\partial_{\mathbf{v}_{0}}]$ applied to $\mathbf{y}$. The second generalized stress $\mathcal{H}$ is a third-order tensor field on $\mathbb{R}^{3\times 2\times 2}$ given by
$\mathcal{H}(\mathbf{y},\theta):=\mathcal{H}_{2}(\mathbf{y},\theta)+2d_{1}L_{\Omega}^{2}\nabla\nabla\mathbf{y},$ (34)
where $\mathcal{H}_{2}$ is obtained by differentiating $W_{2}$ with respect to the Hessian $\nabla\nabla\mathbf{y}$. This differentiation gives
$\mathcal{H}_{2}(\mathbf{y},\theta):=\frac{2c_{2}L_{\Omega}^{2}k(\mathbf{y},\theta)}{|\tilde{\mathbf{u}}_{0}|^{4}|\tilde{\mathbf{v}}_{0}|^{4}}\Big{(}[\mathbf{v}(\theta)\cdot\mathbf{v}^{\prime}(\theta)]\big{\{}\mathbf{n}(\mathbf{y})\otimes\tilde{\mathbf{u}}_{0}\otimes\tilde{\mathbf{u}}_{0}\big{\}}+[\mathbf{u}(\theta)\cdot\mathbf{u}^{\prime}(\theta)]\big{\{}\mathbf{n}(\mathbf{y})\otimes\tilde{\mathbf{v}}_{0}\otimes\tilde{\mathbf{v}}_{0}\big{\}}\Big{)}.$ (35)
The third generalized stress $q$ is a scalar field of the form
$q(\mathbf{y},\theta):=q_{1}(\mathbf{y},\theta)+q_{2}(\mathbf{y},\theta)+2d_{3}\theta$ (36)
with terms $q_{1,2}$ obtained as derivatives of $W_{1,2}$ with respect to $\theta$ via
$q_{1}(\mathbf{y},\theta):=-\frac{2c_{1}}{\sqrt{\det\mathbf{I}(\mathbf{y})}}\Big{(}\mathbf{I}(\mathbf{y})-\mathbf{A}^{T}(\theta)\mathbf{A}(\theta)\Big{)}\colon\big{(}\mathbf{A}^{T}(\theta)\mathbf{A}(\theta)\big{)}^{\prime},$ (37)
$q_{2}(\mathbf{y},\theta):=\frac{2c_{2}L_{\Omega}^{2}k(\mathbf{y},\theta)}{|\tilde{\mathbf{u}}_{0}|^{4}|\tilde{\mathbf{v}}_{0}|^{4}}\Big{(}\big{[}\mathbf{v}(\theta)\cdot\mathbf{v}^{\prime}(\theta)\big{]}^{\prime}\big{[}\tilde{\mathbf{u}}_{0}\cdot\mathbf{II}(\mathbf{y})\tilde{\mathbf{u}}_{0}\big{]}+\big{[}\mathbf{u}(\theta)\cdot\mathbf{u}^{\prime}(\theta)\big{]}^{\prime}\big{[}\tilde{\mathbf{v}}_{0}\cdot\mathbf{II}(\mathbf{y})\tilde{\mathbf{v}}_{0}\big{]}\Big{)}.$
The last generalized stress $\mathbf{j}(\theta)$ in (31) is
$\mathbf{j}(\theta):=2d_{2}L_{\Omega}^{2}\nabla\theta.$ (38)
This completes our description of the weak formulation of the equilibrium equations. In summary, if $(\mathbf{y},\theta)\in V_{\Gamma}$ is a local minimizer to the energy $E$, then (31) holds for all $(\mathbf{v},\eta)\in V_{0}$ for the generalized stresses $\mathbf{P},\mathcal{H},q,\mathbf{j}$ defined by (32-38).
We now derive the strong form of the equilibrium equations and the corresponding natural boundary conditions under the assumption that the minimizers $(\mathbf{y},\theta)$ are sufficiently smooth. A key step in this result is an identity (LABEL:eq:importForStrongForm) on the first integral in (31) derived in Appendix B. We manipulate the weak form using this identity to read
00 0 $=\int_{\Omega}\big{(}\operatorname{div}\big{[}\operatorname{div}\mathcal{H}(\mathbf{y},\theta)-\mathbf{P}(\mathbf{y},\theta)\big{]}-\overline{\mathbf{b}}\big{)}\cdot\mathbf{w}\mathop{}\!\mathrm{d}x+\int_{\Gamma_{\text{t}}}\big{(}\big{[}\mathbf{P}(\mathbf{y},\theta)-\nabla\mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{I}+\mathbf{n}^{\perp}\otimes\mathbf{n}^{\perp})\big{]}\mathbf{n}-\overline{\mathbf{t}}\big{)}\cdot\mathbf{w}\mathop{}\!\mathrm{d}s$ (39)
$\qquad+\int_{\Gamma_{\text{m}}}\big{[}\mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{n}\otimes\mathbf{n})-\overline{\mathbf{m}}\big{]}\cdot(\nabla\mathbf{w})\mathbf{n}\mathop{}\!\mathrm{d}s+\int_{\Omega}\big{(}q(\mathbf{y},\theta)-\nabla\cdot\mathbf{j}(\theta)\big{)}\eta\mathop{}\!\mathrm{d}x+\int_{\partial\Omega}\big{(}\mathbf{j}(\theta)\cdot\mathbf{n}\big{)}\eta\mathop{}\!\mathrm{d}s$
for all $(\mathbf{w},\eta)\in V_{0}$, after employing that $\mathbf{w}=\mathbf{0}$ a.e. on $\Gamma_{\text{d}}$ and $(\nabla\mathbf{w})\mathbf{n}=\mathbf{0}$ a.e. on $\Gamma_{\text{n}}$ to simplify the boundary terms. For clarity on this formula, $\mathbf{n}$ denotes the outward normal to $\partial\Omega$, with $\mathbf{n}^{\perp}=\mathbf{R}(\pi/2)\mathbf{n}$ the corresponding unit tangent vector. In addition, $[\operatorname{div}\mathcal{H}]_{i\alpha}=[\mathcal{H}]_{i\alpha\beta,\beta}$, $[\mathcal{H}\colon(\mathbf{n}\otimes\mathbf{n})]_{i}=[\mathcal{H}]_{i\alpha\beta}n_{\alpha}n_{\beta}$, and $[\nabla\mathcal{H}\colon(\mathbf{I}-\mathbf{n}^{\perp}\otimes\mathbf{n}^{\perp})]_{i\alpha}=[\mathcal{H}]_{i\alpha\beta,\gamma}(\delta_{\beta\gamma}-n^{\perp}_{\beta}n^{\perp}_{\gamma})$. To complete the derivation, we test (39) against various subsets of $(\mathbf{w},\eta)\in V_{0}$ to isolate different terms in the integral and apply standard localization arguments case-by-case to deduce the full set of equilibrium equations and natural boundary conditions. Skipping the details, the governing equations in our constitutive model for parallelogram origami are
$\begin{cases}\operatorname{div}\big{[}\operatorname{div}\mathcal{H}(\mathbf{y},\theta)-\mathbf{P}(\mathbf{y},\theta)\big{]}=\overline{\mathbf{b}}&\text{ in $\Omega$}\\ \nabla\cdot\mathbf{j}(\theta)-q(\mathbf{y},\theta)=0&\text{ in }\Omega\\ \mathbf{y}=\overline{\mathbf{y}}&\text{ on }\Gamma_{\text{d}}\\ \big{[}\mathbf{P}(\mathbf{y},\theta)-\nabla\mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{I}+\mathbf{n}^{\perp}\otimes\mathbf{n}^{\perp})\big{]}\mathbf{n}=\overline{\mathbf{t}}&\text{ on }\Gamma_{\text{t}}\\ (\nabla\mathbf{y})\mathbf{n}=\overline{\mathbf{s}}&\text{ on }\Gamma_{\text{n}}\\ \mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{n}\otimes\mathbf{n})=\overline{\mathbf{m}}&\text{ on }\Gamma_{\text{m}}\\ \nabla\theta\cdot\mathbf{n}=0&\text{ on }\partial\Omega.\end{cases}$ (40)
The complexity of these equations underlies the fact that a continuum description of metamaterials requires not only a characterization of the cell-averaged effective deformation $\mathbf{y}$ but also auxiliary field(s) (in this case an angle field $\theta$) that track the mechanical behavior of the cell-by-cell microstructure. Such generalized elastic continuum are known to be quite rich; see for instance Eringen’s work on microcontinuum theories [21]. The first equilibrium equation is the standard one in non-linear elasticity where $\mathbf{P}(\mathbf{y},\theta)-\operatorname{div}\mathcal{H}(\mathbf{y},\theta)$ plays the role of a generalized stress. Indeed, by freezing $\theta$ (so it is no longer an elastic field variable) and dropping the terms in the model depending on $\nabla\nabla\mathbf{y}$, the governing equations in (40) reduce to the familiar ones: $\operatorname{div}\mathbf{P}(\mathbf{y})+\overline{\mathbf{b}}=\mathbf{0}$ in $\Omega$ subject to mixed traction and Dirichlet boundary conditions $\mathbf{P}(\mathbf{y})\mathbf{n}=\overline{\mathbf{t}}$ on $\Gamma_{\text{t}}$ and $\mathbf{y}=\overline{\mathbf{y}}$ on $\Gamma_{\text{d}}$. Of course, $\theta$ is not frozen but instead an elastic field variable reflecting the actuation of the folds, which helps to reduce the stored elastic energy and stress of the origami under loads. This necessitates another equilibrium equation $\nabla\cdot\mathbf{j}(\theta)-q(\mathbf{y},\theta)=0$ in $\Omega$ and boundary condition $\nabla\theta\cdot\mathbf{n}=0$ on $\partial\Omega$. The higher order gradients of the deformation in the model account for origami bending. They introduce interesting couplings in the equilibrium equations through the terms $\mathcal{H}(\mathbf{y},\theta)$, $\mathbf{P}(\mathbf{y},\theta)$ and $q(\mathbf{y},\theta)$ and lead to somewhat peculiar looking traction and generalized moment boundary conditions. Such boundary conditions, while not an everyday occurrence in modern continuum mechanics research, are completely consistent with those studied by Toupin [60], back when the development of generalized continuum and strain gradient models was a topic of the times.
4 Finite element formulation
In this section, we develop a finite element formulation of the governing equations in (40). The main technical challenge is to address the second gradients of deformation, present in the stored energy of the model, in a computationally reasonable way. Traditional finite element discretizations of the deformation, involving Lagrange polynomials, possess jumps in their derivatives at internal edges and thus are non-conforming in $H^{2}(\Omega,\mathbb{R}^{3})$. This means that such discretizations do not accurately capture the $\nabla\nabla\mathbf{y}$ terms present in our model in a straightforward manner. While $H^{2}$-conforming finite elements have been known for a long time [3, 6], ensuring the desired smoothness across the edges leads, in practice, to elements of order at least $\geq 4^{\text{th}}$, which become computationally expensive. Writing an interpolator for these elements is also difficult, meaning that one often ends up imposing the Dirichlet boundary conditions weakly. Another approach is to use subdivision elements [10], which achieve $H^{2}$-conforming deformations through meshing strategies that originate from computational geometry [34]. This approach has proven successful at simulating the large deformation response of plates and shells, but is difficult to implement in standard finite element packages. We follow instead the interior penalty method in [20]. Specifically, we use traditional Lagrange polynomials for the discretization, allowing the normal derivatives to jump across edges, but then modify the weak formulation of the governing equations to handle these jumps.
4.1 Preliminaries
Here we briefly collect some notation helpful for concisely describing the finite element formulation of (40) using the interior penalty method, which we develop in the next section.
We always consider a discretization of the domain $\Omega$ via a shape-regular conforming triangulation $\mathcal{T}^{h}$ with $h_{T}:=\text{diam}(T)$ and $h:=\max_{T\in\mathcal{T}_{h}}h_{T}$ (see [9] for details). We denote by $\mathcal{E}^{h}$ the set of all edges of $\mathcal{T}^{h}$, $\mathcal{E}_{\Gamma}^{h}$ the set of all boundary edges $e\in\mathcal{E}^{h}$ such that $e\subset\partial\Omega$, and $\mathcal{E}_{\text{int}}^{h}:=\mathcal{E}^{h}\setminus\mathcal{E}_{\Gamma}^{h}$ the set of all internal edges. We also consider the subsets $\mathcal{E}_{\Gamma}^{h}$ denoted $\mathcal{E}_{\Gamma_{\text{d}}}^{h},\mathcal{E}_{\Gamma_{\text{t}}}^{h},\mathcal{E}_{\Gamma_{\text{n}}}^{h},\mathcal{E}_{\Gamma_{\text{m}}}^{h}$, which collect the edges on $\Gamma_{\text{d}}$, $\Gamma_{\text{t}},$ $\Gamma_{\text{n}}$ and $\Gamma_{\text{m}}$, respectively. For any interior edge $e\in\mathcal{E}_{\text{int}}^{h}$ shared by two triangles, denoted $T^{+}$ and $T^{-}$, we define $\mathbf{n}_{e}$ as the unit normal to $e$ that points from $T^{+}$ to $T^{-}$. As field variables defined on the triangulation $\mathcal{T}^{h}$ can jump across edges, we define the jump and average of a field variable ()($\cdot$)( ⋅ ) (any scalar, vector, or $n^{\text{th}}$-order tensor field) across the edge $e$ by $\llbracket(\cdot)\rrbracket_{e}:=(\cdot)^{+}-(\cdot)^{-}$ and {{()}}$\{\mskip-5.0mu\{(\cdot)\}\mskip-5.0mu\}_{e}:=\frac{1}{2}((\cdot)^{+}+(\cdot)^{-})$, respectively, where $(\cdot)^{\pm}$ denotes the limit of the restriction of the field variable $(\cdot)|_{T^{\pm}}$ to the edge $e$. For an exterior edge $e\in\mathcal{E}_{\Gamma}^{h}$, we take $\mathbf{n}_{e}\equiv\mathbf{n}$ to be the outward unit normal of $T$ on $e$. Finally, the length of any edge $e\in\mathcal{E}^{h}$ is given by $|e|$.
We define the approximation space for the deformation and angle field on our triangulation $\mathcal{T}^{h}$ as
$V^{h}:=\big{\{}(\mathbf{y}_{h},\theta_{h})\in C^{0}(\Omega,\mathbb{R}^{3})\times C^{0}(\Omega,\mathbb{R})\colon\mathbf{y}_{h}|_{T}\in P^{2}(T)\;\;\text{and}\;\;\theta_{h}|_{T}\in P^{1}(T)\;\;\forall\;\;T\in\mathcal{T}^{h}\big{\}}$ (41)
where $P^{k}(T)$ denotes the space of Lagrange polynomials of degree $\leq k$ on the triangle $T$, and $C^{0}$ denotes the space of continuous functions. Note that $\mathbf{y}_{h}$ need not be in $H^{2}$, which makes the proposed method nonconforming. We write the classical Scott–Zhang interpolator [22] for $V^{h}$ as $\mathcal{I}_{h}$ and take into account Dirichlet boundary conditions on $\mathbf{y}_{h}$ by defining
$V^{h}_{\Gamma_{\text{d}}}:=\{(\mathbf{y}_{h},\theta_{h})\in V^{h}\colon\mathbf{y}_{h}=\mathcal{I}_{h}\overline{\mathbf{y}}\text{ on }\Gamma_{\text{d}}\}.$ (42)
The Dirichlet boundary condition on the slope, $\nabla\mathbf{y}_{h}\mathbf{n}=\overline{\mathbf{s}}$ on $\Gamma_{\text{n}}$, are imposed weakly. The discrete solution space is then $V^{h}_{\Gamma_{\text{d}}}$. The associated homogeneous space is defined as $V_{0}^{h}$, and is given by (42) but with $\overline{\mathbf{y}}=\mathbf{0}$. A generic deformation $\mathbf{y}_{h}$ in $V^{h}$ can have a discontinuous gradient across $e\in\mathcal{E}^{h}_{\text{int}}$, so we cannot directly define a Hessian (though we can define a gradient in the usual sense as $\mathbf{y}_{h}$ is continuous across $e$). We define instead a broken Hessian $\nabla_{h}\nabla\mathbf{y}_{h}:=\sum_{T\in\mathcal{T}_{h}}\nabla(\nabla\mathbf{y}_{h})_{T}\chi_{T}$, where $\chi_{T}$ is the indicator function of $T\in\mathcal{T}^{h}$, which consists of computing the Hessian elementwise (see [16] for more details). Finally, for any $(\mathbf{y}_{h},\theta_{h})\in V^{h}$, we write the generalized stresses via a slight abuse of notation as $\mathbf{P}(\mathbf{y}_{h},\theta_{h})$, $\mathcal{H}(\mathbf{y}_{h},\theta_{h})$, $q(\mathbf{y}_{h},\theta_{h})$, where terms that depend on the Hessian of $\mathbf{y}_{h}$ (which is not well-defined) are replaced by the broken Hessian $\nabla_{h}\nabla\mathbf{y}_{h}$.
4.2 Weak formulation using the interior penalty method
The problem of solving for an equilibrium solution of the governing equations in (40) can be written succinctly in weak form as follows: Find a $(\mathbf{y},\theta)\in V_{\Gamma}$ such that
$\mathscr{A}_{0}((\mathbf{y},\theta);(\mathbf{w},\eta))=\mathscr{B}_{0}(\mathbf{w})\quad\text{ for all }(\mathbf{w},\theta)\in V_{0}$ (43)
where, following (31), $\mathscr{A}_{0}((\mathbf{y},\theta);(\mathbf{w},\eta)):=\int_{\Omega}\{\mathbf{P}(\mathbf{y},\theta)\colon\nabla\mathbf{w}+\big{\langle}\mathcal{H}(\mathbf{y},\theta),\nabla\nabla\mathbf{w}\big{\rangle}+q(\mathbf{y},\theta)\eta+\mathbf{j}(\theta)\cdot\nabla\eta\}\mathop{}\!\mathrm{d}x$ and $\mathscr{B}_{0}(\mathbf{w}):=\int_{\Omega}\overline{\mathbf{b}}\cdot\mathbf{w}\mathop{}\!\mathrm{d}x+\int_{\Gamma_{\text{t}}}\overline{\mathbf{t}}\cdot\mathbf{w}\mathop{}\!\mathrm{d}s+\int_{\Gamma_{\text{m}}}\overline{\mathbf{m}}\cdot(\nabla\mathbf{w})\mathbf{n}\mathop{}\!\mathrm{d}s$ denote the parts of the weak form associated to the stored elastic energy and external loads, respectively. Of course, solving for the equilibrium solution numerically necessitates some approximations.
We now formulate conditions for an approximate equilibrium solution using $C^{0}$ finite elements and the interior penalty method. In short, after discretizing the domain, deformation and angle field via the description in the previous section, we modify the weak formulation in (43) to read: Find a $(\mathbf{y}_{h},\theta_{h})\in V^{h}_{\Gamma_{\text{d}}}$ such that
$\mathscr{A}^{h}((\mathbf{y}_{h},\theta_{h});(\mathbf{w}_{h},\eta_{h}))=\mathscr{B}_{0}(\mathbf{w}_{h})\quad\text{ for all }(\mathbf{w}_{h},\theta_{h})\in V^{h}_{0}$ (44)
The left-hand side of this equation is the sum of four terms
$\mathscr{A}^{h}:=\mathscr{A}^{h}_{0}+\mathscr{A}^{h}_{\text{con}}+\mathscr{A}^{h}_{\text{sta}}+\mathscr{A}^{h}_{\text{bnd}},$ (45)
each evaluated on $(\mathbf{y}_{h},\theta_{h})\in V^{h}_{\Gamma_{\text{d}}}$ and $(\mathbf{w}_{h},\eta_{h})\in V^{h}_{0}$. The first term is a discrete analog of the internal energy term $\mathscr{A}_{0}((\mathbf{y},\theta);(\mathbf{w},\eta))$ in (43). It is given by
$\mathscr{A}^{h}_{0}((\mathbf{y}_{h},\theta_{h});(\mathbf{w}_{h},\eta_{h})):=\sum_{T\in\mathcal{T}^{h}}\int_{T}\Big{\{}\mathbf{P}(\mathbf{y}_{h},\theta_{h})\colon\nabla\mathbf{w}_{h}+\big{\langle}\mathcal{H}(\mathbf{y}_{h},\theta_{h}),\nabla_{h}\nabla\mathbf{w}_{h}\big{\rangle}+q(\mathbf{y}_{h},\theta_{h})\eta_{h}+\mathbf{j}(\theta_{h})\cdot\nabla\eta_{h}\Big{\}}\mathop{}\!\mathrm{d}x.$ (46)
The second term $\mathscr{A}_{\text{con}}^{h}$ is introduced for consistency, i.e., to ensure that any equilibrium solution to the original problem in (40) is also a solution to the weak form in our numerical method. It is of the form
$\mathscr{A}^{h}_{\text{con}}((\mathbf{y}_{h},\theta_{h});(\mathbf{w}_{h},\eta_{h})):=-\sum_{e\in\mathcal{E}_{\text{int}}^{h}}\int_{e}\big{(}\{\mskip-5.0mu\{\mathcal{H}(\mathbf{y}_{h},\theta_{h})\}\mskip-5.0mu\}_{e}\cdot\mathbf{n}_{e}\big{)}\colon\llbracket\nabla\mathbf{w}_{h}\rrbracket_{e}\mathop{}\!\mathrm{d}s-\int_{\Gamma_{\text{n}}}\big{(}\mathcal{H}(\mathbf{y}_{h},\theta_{h})\cdot\mathbf{n}\big{)}\colon\nabla\mathbf{w}_{h}\mathop{}\!\mathrm{d}s,$ (47)
where the integrals over each $e\in\mathcal{E}_{\text{int}}^{h}$ arise due to the jumps in both $\nabla\mathbf{y}_{h}$ and $\nabla\mathbf{w}_{h}$ along edges in the mesh, and the one over $\Gamma_{\text{n}}$ results from imposing the slope boundary conditions weakly in this formulation. Appendix C provides a derivation of (47) (see also [16, 20] for further examples of such terms and their justification). The third term $\mathscr{A}_{\text{sta}}^{h}$ seeks to stabilize/minimize the aforementioned jumps in the deformation gradient via
$\mathscr{A}^{h}_{\text{sta}}((\mathbf{y}_{h},\theta_{h});(\mathbf{w}_{h},\eta_{h})):=\sum_{e\in\mathcal{E}^{h}_{\text{int}}}\frac{\alpha}{|e|}\int_{e}\llbracket\nabla\mathbf{y}_{h}\rrbracket_{e}:\llbracket\nabla\mathbf{w}_{h}\rrbracket_{e}\mathop{}\!\mathrm{d}s.$ (48)
It corresponds to a jump energy of the form $\tfrac{1}{2}\sum_{e\in\mathcal{E}^{h}_{\text{int}}}\frac{\alpha}{|e|}\int_{e}\llbracket_{e}\nabla\mathbf{y}_{h}\rrbracket^{2}\mathop{}\!\mathrm{d}x$ and provides coercivity of the overall discrete energy for $\alpha>0$ large enough, ensuring the existence of a solution to (44). The final term $\mathscr{A}_{\text{bnd}}^{h}$ seeks to impose the slope boundary condition weakly via
$\mathscr{A}^{h}_{\text{bnd}}((\mathbf{y}_{h},\theta_{h});(\mathbf{w}_{h},\eta_{h})):=\sum_{e\in\mathcal{E}_{\Gamma_{\text{n}}}^{h}}\frac{\alpha}{|e|}\int_{e}\big{(}(\nabla\mathbf{y}_{h})\mathbf{n}-\bar{\mathbf{s}}\big{)}\cdot(\nabla\mathbf{w}_{h})\mathbf{n}\mathop{}\!\mathrm{d}s.$ (49)
It arises from an energy of the form $\tfrac{1}{2}\sum_{e\in\mathcal{E}_{\Gamma_{\text{n}}}^{h}}\tfrac{\alpha}{|e|}\int_{e}|(\nabla\mathbf{y}_{h})\mathbf{n}-\bar{\mathbf{s}}|^{2}\mathop{}\!\mathrm{d}s$, and thus furnishes $(\nabla\mathbf{y}_{h})\mathbf{n}\approx\overline{\mathbf{s}}$ by way of a least square fit.
4.3 Implementation
The weak form in (44) is a nonlinear algebraic system of equations. We implement it in Firedrake [29], which is a finite element library with a Python interface that allows one to choose a finite element space (41) and then write the discrete weak formulation (44). We use Firedrake’s integrated Newton solver to find a solution to (44). As the convergence of Newton solvers are sensitive to the initial guess, we provide a method to compute an initial guess for $(\mathbf{y}_{h},\theta_{h})$ in Appendix D. Also, we typically employ a loading protocol that monotonically increases the load in small increments until finally solving the desired large deformation boundary value problem. We use the method in Appendix D to initialize the simulation in the first step, while all subsequent steps are initialized using the converged solution of the prior increment. Finally, it is well-known that penalty methods, like the one proposed here, are sensitive to the choice of $\alpha$. By trial and error, we have found that $\alpha=0.1$ works well when $c_{1}\in[1,5]$ and the other moduli $c_{2},d_{1},\ldots,d_{3}$ are consistent with the scalings discussed in Section 3.1. We use this value for $\alpha$ in all forthcoming simulations in Section 5.
5 Examples
This section showcases a variety of examples of parallelogram origami under loads, modeled using the constitutive framework in Section 3 along with the numerical method in Section 4. We explore the effective mechanical behavior of the two canonical examples of parallelogram origami: Miura and Eggbox origami. Miura is an auxetic metamaterial with a negative Poisson’s ratio; Eggbox is not auxetic and thus has a positive Poisson’s ratio. Both patterns posses unit cells built from the single parallelogram panel and have orthogonal Bravais lattice vectors $\mathbf{u}(\theta)=\lambda_{1}(\theta)\mathbf{e}_{1}$ and $\mathbf{v}(\theta)=\lambda_{2}(\theta)\mathbf{e}_{2}$ with simple trigonometric expressions, making them convenient for analytical and numerical investigation. We focus on examples where the unit cell is composed of four identical rhombi panels of side length 1 and acute angle $\pi/3$ (see Fig. 5(a) and (d)). We also take the reference configuration of all the examples in this section to consist of the same partially folded unit cells (Fig. 5(b) for Miura origami, Fig. 5(e) for Eggbox). After some tedious algebra (which we do not detail), the actuation of the Bravais lattice vectors for Miura origami in this setting is given by stretches $\lambda_{1}(\theta)\equiv\lambda_{1}^{(\text{MO})}(\theta)$ and $\lambda_{2}(\theta)\equiv\lambda_{2}^{(\text{MO})}(\theta)$ that satisfy
$\lambda_{1}^{(\text{MO})}(\theta)=\sqrt{3}\cos\left(\frac{\theta+\pi/6}{2}\right),\quad\lambda_{2}^{(\text{MO})}(\theta)=2\sqrt{2}\Big{(}5-3\cos(\theta+\pi/6)\Big{)}^{-1/2},\quad\theta\in\left(-\frac{\pi}{6},\frac{5\pi}{6}\right).$ (50)
For Eggbox origami, the analogous stretches $\lambda_{1}(\theta)\equiv\lambda_{1}^{(\text{EO})}(\theta)$ and $\lambda_{2}(\theta)\equiv\lambda_{2}^{\text{(EO)}}(\theta)$ are
$\lambda_{1}^{(\text{EO})}(\theta)=2\sin\Big{[}\frac{1}{2}\Big{(}\arccos\big{(}1-\cos\theta\big{)}-\theta\Big{)}\Big{]},\quad\lambda_{2}^{(\text{EO})}(\theta)=\lambda_{1}^{(\text{EO})}(-\theta),\quad\theta\in\big{(}-\frac{\pi}{3},\frac{\pi}{3}\big{)}.$ (51)
Fig. 5(c) and (f) illustrate the actuation. In all examples, whether Miura or Eggbox, we take the effective reference domain $\Omega$ to be
$\Omega=(0,\lambda_{1}(0))\times(0,\lambda_{2}(0))\approx\begin{cases}(0,1.67)\times(0,1.82)&\text{ for Miura}\\ (0,1.41)\times(0,1.41)&\text{ for Eggbox}.\end{cases}$ (52)
This choice, while seemingly more cumbersome than, e.g., fixing a unit square domain, allows us to construct origami deformations of $M\times M$ cells that match the analytical and simulation results of the effective model.
Note that the design space of all parallelogram origami is large (in fact, 7 dimensional [^6] A parallelogram origami unit cell has 4 lengths and 4 sector angles corresponding to the sides of the parallelograms and their shape. However, one of the lengths can be fixed (set to 1, for instance) without loss of generality since the elastic model does not change under a dilation of the pattern.), yet we are focusing on just two designs here. This is purposeful. Our primary interest in this work is to introduce and validate the model and numerical method. Due to their familiarity, the Miura and Eggbox patterns make excellent templates for this validation. Also, as we will show, these patterns turn out to be quite rich in their own right, with many distinguishing features worthy of a thorough investigation and comparison. A more detailed examination of the behavior of broad classes of parallelogram origami under loads will be the topic of future work.
Figure 5: The Miura and Eggbox origami unit cells used in our catalog of examples. (a,d) Flat states. Each unit cell is built using four rhombi panels with side lengths 1 and acute angle $\pi/3$. (b,e) Partially folded reference states. $\lambda_{1}(0)$ and $\lambda_{2}(0)$ denote the side lengths from the formulas in (50) and (51), evaluated at $\theta=0$. (c,f) Actuated states. The side lengths change from $\lambda_{i}(0)$ to $\lambda_{i}(\theta)$, $i=1,2$, under the actuation, which is parameterized by $\theta$.
5.1 Mechanism deformations
We first study the effective mechanical response of parallelogram origami patterns under a pure mechanism actuation. This setting corresponds to Dirichlet boundary conditions of the form $\overline{\mathbf{y}}(\mathbf{x})=\mathbf{A}(\overline{\theta})\mathbf{x}$ on $\partial\Omega$, where $\mathbf{A}(\theta)$ is the shape tensor in (8) and $\overline{\theta}\in(\theta^{-},\theta^{+})$ sets the level of actuation. We also assume that $\overline{\mathbf{b}}=\overline{\mathbf{m}}=\mathbf{0}$, so there are no applied forces. Thus, in the absence of a mechanical resistance to folding ($d_{3}=0$), the energy minimizing (and zero energy) deformation and angle field in the model are that of the purely mechanistic response
$\mathbf{y}(\mathbf{x})=\mathbf{A}(\overline{\theta})\mathbf{x},\quad\theta(\mathbf{x})=\overline{\theta}\quad\text{ in }\Omega.$ (53)
The situation is more complicated when $d_{3}>0$. The folds prefer not to be actuated $(\theta=0)$, leading to stresses that depend non-linearly on $\overline{\theta}$.
Figure 6: Normalized force versus compressive strain for the pure mechanism in Miura and Eggbox origami. (a) The force plot for Miura origami, and (b) its representative actuation for six points (i)-(vi) along the curve. (c) The force plot for Eggbox origami, and (d) its representative actuation for six points (I)-(VI) along the curve. The normalized force and compressive strain are defined in (LABEL:eq:normForce).
We obtain an analytical expression for the stresses as follows. First, observe that the ansatz $(\mathbf{y}(\mathbf{x}),\theta(\mathbf{x}))=(\mathbf{A}(\overline{\theta})\mathbf{x},\overline{\theta}+\delta\theta)$ resolves the boundary conditions and the first equilibrium equation in (40) automatically. The second equilibrium equation $\nabla\cdot\mathbf{j}(\theta)-q(\mathbf{y},\theta)=0$ reduces to a purely algebraic equation of the form
$q_{1}(\mathbf{A}(\overline{\theta})\mathbf{x},\overline{\theta}+\delta\theta)+2d_{3}(\overline{\theta}+\delta\theta)=0$ (54)
for $q_{1}(\mathbf{y},\theta)$ in (36). Since $q_{1}(\mathbf{A}(\overline{\theta})\mathbf{x},\bar{\theta})=0$ and $\partial_{\theta}q_{1}(\mathbf{A}(\overline{\theta})\mathbf{x},\bar{\theta})\neq 0$, this equation has a solution $\delta\theta$ provided $d_{3}/c_{1}>0$ is sufficiently small, which is the physically relevant regime. To approximate the solution, note that (54) satisfies
$\frac{3}{2}g^{\prime}(\overline{\theta})\delta\theta^{2}+2\big{(}g(\overline{\theta})+j(\overline{\theta})\tfrac{d_{3}}{c_{1}}\big{)}\delta\theta+2\frac{d_{3}}{c_{1}}j(\overline{\theta})\overline{\theta}=O(\delta\theta^{3})$ (55)
for $g(\overline{\theta}):=|\big{(}\mathbf{A}^{T}(\overline{\theta})\mathbf{A}(\overline{\theta})\big{)}^{\prime}|^{2}$ and $j(\overline{\theta}):=\sqrt{\det\big{(}\mathbf{A}^{T}(\overline{\theta})\mathbf{A}(\overline{\theta})\big{)}}$. Hence, neglecting the terms of order $O(\delta\theta^{3})$ leads to
$\delta\theta\approx-2\frac{\tfrac{d_{3}}{c_{1}}j(\overline{\theta})\overline{\theta}}{\big{(}g(\overline{\theta})+j(\overline{\theta})\tfrac{d_{3}}{c_{1}}\big{)}}\left(1+\sqrt{1-3\frac{\tfrac{d_{3}}{c_{1}}g^{\prime}(\overline{\theta})j(\overline{\theta})\overline{\theta}}{\big{(}g(\overline{\theta})+j(\overline{\theta})\tfrac{d_{3}}{c_{1}}\big{)}^{2}}}\right)^{-1}$ (56)
using the quadratic formula. The Piola-Kirchhoff stress at quadratic order is then
$\mathbf{P}\big{(}\mathbf{A}(\overline{\theta})\mathbf{x},\overline{\theta}+\delta\theta\big{)}\approx-4\frac{c_{1}}{j(\overline{\theta})}\mathbf{A}(\overline{\theta})\Big{[}\delta\theta\big{(}\mathbf{A}^{T}(\overline{\theta})\mathbf{A}(\overline{\theta})\big{)}^{\prime}+\tfrac{\delta\theta^{2}}{2}\big{(}\mathbf{A}^{T}(\overline{\theta})\mathbf{A}(\overline{\theta})\big{)}^{\prime\prime}+\tfrac{\delta\theta^{2}}{4}g(\overline{\theta})\big{(}\mathbf{A}^{T}(\overline{\theta})\mathbf{A}(\overline{\theta})\big{)}^{-1}\Big{]}.$ (57)
As these formulas highlight, even the simplest example of a pure mechanism has a rich nonlinear relationship between mechanical deformation and load. Fig 6(a-b) illustrate the mechanical response concretely in the context of Miura origami; Fig 6(c-d) do likewise for Eggbox origami. Since these examples possess orthogonal Bravais lattices vectors (in (50) and (51)), we plot the normalized compressive force in the $\mathbf{e}_{1}$ direction versus the compressive strain in that direction using the formulas
$f_{1}(\overline{\theta}):=-\lambda_{2}(0)\frac{[\mathbf{P}\big{(}\mathbf{A}(\overline{\theta})\mathbf{x},\overline{\theta}+\delta\theta\big{)}]_{11}}{d_{3}}\quad\text{ and }\quad$ $\varepsilon_{1}(\overline{\theta}):=-\frac{\lambda_{1}(\overline{\theta})-\lambda_{1}(0)}{\lambda_{1}(0)}.$ (58)
Notice from the expressions in (56) and (57) that the folding modulus $d_{3}$, not the bulk modulus $c_{1}$, sets the scale for the force. Furthermore, the expression for the normalized force depends on the ratio $d_{3}/c_{1}$. We therefore consider four decades of $d_{3}/c_{1}$ in the plots in Fig. 6. We also compare the exact solution for the value $d_{3}/c_{1}=2\times 10^{-2}$ to the one obtained using our numerical method, and to linear and quadratic approximations of the curves. [^7] The linear curve is obtained by replacing the exact $\delta\theta$ solving (54) with $\delta\theta_{\text{lin}}=-2\tfrac{d_{3}}{c_{1}}j(\overline{\theta})\overline{\theta}/\big{(}g(\overline{\theta})+j(\overline{\theta})\tfrac{d_{3}}{c_{1}}\big{)}$ and $\mathbf{P}(\mathbf{A}(\overline{\theta})\mathbf{x},\overline{\theta}+\delta\theta)$ in the formula for $f_{1}(\overline{\theta})$ with $-4\big{(}c_{1}\delta\theta_{\text{lin}}/j(\overline{\theta})\big{)}\mathbf{A}(\overline{\theta})\big{(}\mathbf{A}^{T}(\overline{\theta})\mathbf{A}(\overline{\theta})\big{)}^{\prime}$. The quadratic curve is obtained by using the approximations in (56) and (57). The agreement between the exact and numerical solutions are essentially perfect in both examples throughout the entire actuation process. However, the two approximating curves deviate markedly at a large values of $\overline{\theta}$, presumably because both $g(\overline{\theta})$ and $j(\overline{\theta})$ approach 00 as $\overline{\theta}$ goes to the fully folded states in Miura and Eggbox origami, making the Taylor expansions unreliable at large actuation.
Perhaps the most interesting aspect of these plots is the stress-softening behavior. This type of behavior was first theoretically reported in [63] and subsequently reproduce in a bar and hinge model in [33]. We see from our curves that this loss of monotonicity occurs when $d_{3}/c_{1}$ is sufficiently small. The curves also quickly spring back up, a behavior that arises because the bulk energy $W_{1}(\theta,\mathbf{I}(\mathbf{y}))$ in (20-LABEL:eq:W12Def) is scaled by $1/\sqrt{\det\mathbf{I}(\mathbf{y})}$, which elastically penalizes deformations where the panels are close to overlapping one another. In particular, when the actuation approaches the fully folded states (see Fig. 6(vi) and (VI)), the second term in the stress $\mathbf{P}_{1}(\mathbf{y},\theta)$ in (LABEL:eq:P1P2) dominates the first, leading the force curves to veer upwards.
It is also interesting to compare and contrast the curves for Miura and Eggbox. We see that both sets of curves are qualitatively similar — the normlized force increases sub-linearly, before dipping a bit (depending on the ratio $d_{3}/c_{1})$ and then springing back up. However, most notably, the Miura pattern experiences significantly larger normalized force compared to the Eggbox one, even though the comparison is for the same underlying moduli parameters $c_{1}$ and $d_{3}$. This feature highlights the dominant role of geometry in this calculation. To get more granular, observe that the normalized stiffness of each pattern in its undeformed configuration is given by
$\frac{f_{1}^{\prime}(0)}{\varepsilon_{1}^{\prime}(0)}=\frac{8}{g(0)+\tfrac{d_{3}}{c_{1}}},$ (59)
where $g(0)$ depends only on the initial geometry of the pattern. It happens that $g^{(\text{MO})}(0)\approx 0.461$ for Miura and $g^{(\text{EO})}(0)=2$ for Eggbox. Thus, the initial stiffness for the Miura is more than four times larger than that of the Eggbox pattern solely due to geometry. This feature more-or-less persists throughout the entirety of actuation.
5.2 Pure bending and twisting deformations
Having examined the pure mechanism deformations both analytically and numerically, we now explore pure bending and twisting modes. These deformation modes arise as simple but inhomogeneous solutions to the purely geometric constraints on the first and second fundamental forms in (11) and (15), i.e., they come from pairs $(\mathbf{y},\theta)\colon\Omega\rightarrow\mathbb{R}^{3}\times(\theta^{-},\theta^{+})$ that satisfy
$\mathbf{I}(\mathbf{y})=\mathbf{A}^{T}(\theta)\mathbf{A}(\theta),\quad\big{[}\mathbf{v}(\theta)\cdot\mathbf{v}^{\prime}(\theta)\big{]}\big{[}\tilde{\mathbf{u}}_{0}\cdot\mathbf{II}(\mathbf{y})\tilde{\mathbf{u}}_{0}\big{]}+\big{[}\mathbf{u}(\theta)\cdot\mathbf{u}^{\prime}(\theta)\big{]}\big{[}\tilde{\mathbf{v}}_{0}\cdot\mathbf{II}(\mathbf{y})\tilde{\mathbf{v}}_{0}\big{]}=0.$ (60)
In our prior work [64], we developed a large class of such solutions. We briefly reproduce some of that analysis to help explain the origins and nature of these deformation modes (see Section 5.1 and Appendix C in [64] for more details).
Although we eventually specialize to orthogonal parameterizations for $\mathbf{u}(\theta)$ and $\mathbf{v}(\theta)$ (see (50-51)) and rectangular domains (see (52)), we start fairly general here by assuming only that $\Omega$ is simply connected (has no holes). Then, using a variant on the Gauss and Codazzi-Mainardi compatibility conditions from differential geometry [17], the constraints in (60) can be recast equivalently as a coupled system of algebraic and differential equations: find a set of fields $(\bm{\omega}_{\mathbf{u}_{0}},\bm{\omega}_{\mathbf{v}_{0}},\theta)\colon\Omega\rightarrow\mathbb{R}^{3}\times\mathbb{R}^{3}\times(\theta^{-},\theta^{+})$ such that
$\begin{cases}\bm{\omega}_{\mathbf{u}_{0}}\times\mathbf{v}(\theta)+(\partial_{\mathbf{u}_{0}}\theta)\mathbf{v}^{\prime}(\theta)=\bm{\omega}_{\mathbf{v}_{0}}\times\mathbf{u}(\theta)+(\partial_{\mathbf{v}_{0}}\theta)\mathbf{u}^{\prime}(\theta)\\ \bm{\omega}_{\mathbf{u}_{0}}\cdot\mathbf{v}^{\prime}(\theta)=\bm{\omega}_{\mathbf{v}_{0}}\cdot\mathbf{u}^{\prime}(\theta)\\ \partial_{\mathbf{u}_{0}}\bm{\omega}_{\mathbf{v}_{0}}-\partial_{\mathbf{v}_{0}}\bm{\omega}_{\mathbf{u}_{0}}=\bm{\omega}_{\mathbf{v}_{0}}\times\bm{\omega}_{\mathbf{u}_{0}}.\end{cases}$ (61)
To explain the equivalence, there is first, for any solution to (61), a unique rotation field $\mathbf{R}\colon\Omega\rightarrow SO(3)$ that solves the differential equations $\partial_{\mathbf{u}_{0}}\mathbf{R}=\mathbf{R}(\bm{\omega}_{\mathbf{u}_{0}}\times)$ and $\partial_{\mathbf{v}_{0}}\mathbf{R}=\mathbf{R}(\bm{\omega}_{\mathbf{v}_{0}}\times)$ subject to $\mathbf{R}(\langle\mathbf{x}\rangle)=\mathbf{I}$, where $\langle\mathbf{x}\rangle=\frac{1}{|\Omega|}\int_{\Omega}\mathbf{x}\mathop{}\!\mathrm{d}x$ is the average point inside the domain $\Omega$. Furthermore, there is a unique deformation $\mathbf{y}\colon\Omega\rightarrow\mathbb{R}^{3}$ that solves the differential equations $\partial_{\mathbf{u}_{0}}\mathbf{y}=\mathbf{R}\mathbf{u}(\theta)$ and $\partial_{\mathbf{v}_{0}}\mathbf{y}=\mathbf{R}\mathbf{v}(\theta)$ subject to $\mathbf{y}(\langle\mathbf{x}\rangle)=\langle\mathbf{x}\rangle$. [^8] The specification of $\mathbf{R}$ and $\mathbf{y}$ at a point inside the domain, e.g., the average point $\langle\mathbf{x}\rangle$, guarantees uniqueness of the two fields. Without such specification, the fields are unique only up to a rigid body motion. The pair $(\mathbf{y},\theta)$, constructed as above, satisfies (60); moreover, the procedure to obtain $(\mathbf{y},\theta)$ based on a solution to (61) is exhaustive so long as we concern ourselves with sufficiently smooth solutions to (60).
Actually, the system in (61) can be further refined. Notice that the first two constraints are four linear equations in the six DOFs represented by the vectors $\bm{\omega}_{\mathbf{u}_{0}}$ and $\bm{\omega}_{\mathbf{v}_{0}}$. They are solved by writing
$\bm{\omega}_{\mathbf{u}_{0}}=\tau\mathbf{u}(\theta)+\kappa(\mathbf{u}^{\prime}(\theta)\cdot\mathbf{u}(\theta))\mathbf{v}(\theta)+\Big{(}\frac{(\partial_{\mathbf{u}_{0}}\theta)(\mathbf{v}^{\prime}(\theta)\cdot\mathbf{u}(\theta))-(\partial_{\mathbf{v}_{0}}\theta)(\mathbf{u}^{\prime}(\theta)\cdot\mathbf{u}(\theta))}{\mathbf{e}_{3}\cdot(\mathbf{u}(\theta)\times\mathbf{v}(\theta))}\Big{)}\mathbf{e}_{3},$ (62)
$\bm{\omega}_{\mathbf{v}_{0}}=\kappa(\mathbf{v}^{\prime}(\theta)\cdot\mathbf{v}(\theta))\mathbf{u}(\theta)-\tau\mathbf{v}(\theta)+\Big{(}\frac{(\partial_{\mathbf{u}_{0}}\theta)(\mathbf{v}^{\prime}(\theta)\cdot\mathbf{v}(\theta))-(\partial_{\mathbf{v}_{0}}\theta)(\mathbf{u}^{\prime}(\theta)\cdot\mathbf{v}(\theta))}{\mathbf{e}_{3}\cdot(\mathbf{u}(\theta)\times\mathbf{v}(\theta))}\Big{)}\mathbf{e}_{3},$
where $\kappa,\tau\colon\Omega\rightarrow\mathbb{R}$ are now the independent variables along with $\theta$. Substituting this parameterization into the final equation in (61) produces a system of three nonlinear PDEs in the three unknowns $\kappa,\tau$, and $\theta$ (a lengthy but explicit set of equations (140-142) in Section 5.1 of [64]).
Figure 7: Pure bending and twisting modes of Miura origami. (a) Origami deformation for an analytical solution $(\mathbf{y},\theta)$ of a pure bending mode. (b) Corresponding finite element simulation $(\mathbf{y}_{h},\theta_{h})$ under Dirichlet boundary conditions specified by the analytical solution. The error in (c) the effective displacement and (d) the angle field between the analytical and simulated solutions, plotted on the reference domain. For reference, the characteristic out-of-plane displacement in bending is $\sim 0.2$ and $\max|\theta_{h}|=0.6254$. (e-h) Analogous plots for pure twisting. The characteristic out-of-plane displacement is $\sim 0.4$ and $\max|\theta_{h}|=0.4964$.
Figure 8: Pure bending and twisting modes of Eggbox origami. (a) Origami deformation for an analytical solution $(\mathbf{y},\theta)$ of a pure bending mode. (b) Corresponding finite element simulation $(\mathbf{y}_{h},\theta_{h})$ under Dirichlet boundary conditions specified by the analytical solution. The error in (c) the effective displacement and (d) the angle field between the analytical and simulated solutions, plotted on the reference domain. For reference, the characteristic out-of-plane displacement in bending is $\sim 0.3$ and $\max|\theta_{h}|=0.4217$. (e-h) Analogous plots for pure twist. The characteristic out-of-plane displacement is $\sim 0.5$ and $\max|\theta_{h}|=0.5035$.
A parameterization of all possible solutions $(\kappa,\tau,\theta)$ to the PDE system seems out of reach. However, we can produce some illustrative examples based on a 1D ansatz. In the setting where the lattice vectors are orthogonal $\mathbf{u}(\theta)=\lambda_{1}(\theta)\mathbf{e}_{1}$ and $\mathbf{v}(\theta)=\lambda_{2}(\theta)\mathbf{e}_{2}$, like in the Miura and Eggbox patterns, we set
$\theta(\mathbf{x})=\theta_{\text{1D}}\big{(}\mathbf{x}\cdot\lambda_{2}^{-1}(0)\tilde{\mathbf{e}}_{2}\big{)},\quad\kappa(\mathbf{x})=\kappa_{\text{1D}}\big{(}\mathbf{x}\cdot\lambda_{2}^{-1}(0)\tilde{\mathbf{e}}_{2}\big{)},\quad\tau(\mathbf{x})=\tau_{\text{1D}}\big{(}\mathbf{x}\cdot\lambda_{2}^{-1}(0)\tilde{\mathbf{e}}_{2}\big{)},$ (63)
where $\tilde{\mathbf{e}}_{1},\tilde{\mathbf{e}}_{2}$ are the standard basis vectors on $\mathbb{R}^{2}$. Substituting the ansatz in (63) into $(\kappa,\tau,\theta)$ PDE leads to a coupled set of ODEs in the scalar variable $s:=\mathbf{x}\cdot\lambda_{2}^{-1}(0)\tilde{\mathbf{e}}_{2}$ for the fields $(\theta_{\text{1D}}(s),\kappa_{\text{1D}}(s),\tau_{\text{1D}}(s))$, which we solved in Appendix C [64]. The general solutions are
$\kappa_{\text{1D}}=\frac{c_{\kappa}}{\lambda_{1}(\theta_{\text{1D}})\lambda_{1}^{\prime}(\theta_{\text{1D}})\lambda_{2}^{2}(\theta_{\text{1D}})},\quad\tau_{\text{1D}}=\frac{c_{\tau}}{\lambda^{2}_{1}(\theta_{\text{1D}})}$ (64)
for arbitrary constants $c_{\kappa}$ and $c_{\tau}$, and for $\theta_{\text{1D}}$ solving the nonlinear ODE
$\frac{d}{ds}\Big{[}\frac{\lambda_{1}^{\prime}(\theta_{\text{1D}})}{\lambda_{2}(\theta_{\text{1D}})}\frac{d}{ds}\theta_{\text{1D}}\Big{]}=c_{\tau}^{2}\frac{\lambda_{2}(\theta_{\text{1D}})}{\lambda_{1}^{3}(\theta_{\text{1D}})}+c_{\kappa}^{2}\frac{\lambda_{2}^{\prime}(\theta_{\text{1D}})}{\lambda_{2}^{2}(\theta_{\text{1D}})\lambda^{\prime}_{1}(\theta_{\text{1D}})}$ (65)
subject to the initial conditions $\theta_{\text{1D}}(0)=\overline{\theta}$ and $\theta_{\text{1D}}^{\prime}(0)=\overline{\zeta}$.
In this parameterization, $\kappa_{\text{1D}}$ represents bending while $\tau_{\text{1D}}$ represents twist. As always, $\theta_{\text{1D}}$ represents cell-by-cell actuation. Thus, pure bending modes are obtained by solving (64) and (65) with $c_{\tau}=0$; pure twisting modes are obtained likewise by solving these equations with $c_{\kappa}=0$.
Fig. 7 and 8 compare analytical solutions of pure bending and twisting modes for both Miura and Eggbox to analogous simulations using our model and numerical method. In all cases, the $\theta$-dependence is specified by (50) for Miura and (51) for Eggbox and $\Omega$ is as in (52). Concerning the analytics on Miura, the inputs to the ODE in (65) are $(c_{\kappa},c_{\tau},\theta_{\text{1D}}(0),\theta_{\text{1D}}^{\prime}(0))=(1.518,0,-\pi/6+0.2,16)$ for pure bending and $=(0,1.5205,-\pi/6+0.2,11)$ for pure twisting; the analogous inputs for Eggbox are $=(1.43,0,0.4,-1.5)$ and $=(0,1.9,-0.5,4)$ for bend and twist, respectively. These inputs are chosen so that the solution $\theta_{\text{1D}}(s)$ is approximately symmetric about $s=1/2$, corresponding to the midline of the domain $\Omega$ in the $\mathbf{v}_{0}$ direction. After obtaining the solutions, we substitute $(\theta_{\text{1D}},\kappa_{\text{1D}},\tau_{\text{1D}})\equiv(\theta,\kappa,\tau)$ into (LABEL:eq:theOmegaParams). The rotation field $\mathbf{R}$ and deformation $\mathbf{y}$ are then determined by solving the differential equations $\partial_{\mathbf{u}_{0}}\mathbf{R}=\mathbf{R}(\bm{\omega}_{\mathbf{u}_{0}}\times)$, $\partial_{\mathbf{v}_{0}}\mathbf{R}=\mathbf{R}(\bm{\omega}_{\mathbf{v}_{0}}\times)$, $\partial_{\mathbf{u}_{0}}\mathbf{y}=\mathbf{R}\mathbf{u}(\theta),$ and $\partial_{\mathbf{v}_{0}}\mathbf{y}=\mathbf{R}\mathbf{v}(\theta)$, as explained in Appendix C of [64]. Finally, Figs. 7, 8(a) and (e) show plots of origami deformations constructed from these solutions using an ansatz outlined in Appendix F. Notice that the Miura takes the shape of a saddle for pure bending $(c_{\tau}=0)$, while Eggbox takes that of the opposite bending profile in a cap. In contrast, the pure twist deformations of these pattern $(c_{\kappa}=0)$ are qualitatively similar.
Our goal with the numerics is to approximate the above analytical solutions $(\mathbf{y},\theta)$ via corresponding finite element simulations $(\mathbf{y}_{h},\theta_{h})$ as a means to validate the nonlinear parts of our model and numerical implementation. To do this, we impose full Dirichlet boundary conditions $\mathbf{y}_{h}=\mathbf{y}$ and $(\nabla\mathbf{y}_{h})\mathbf{n}=(\nabla\mathbf{y})\mathbf{n}$ on $\partial\Omega$. We also set the moduli parameters of the model for all numerical simulations as $c_{1}=1,c_{2}L_{\Omega}^{2}=0.5,d_{1}L_{\Omega}^{2}=10^{-3},d_{2}=d_{3}=0$. By these choices, the stored elastic energy for $(\mathbf{y}_{h},\theta_{h})$ is dominated mostly by deviations from the purely geometric constraints in (60). In particular, the regularizing terms penalizing the actuation field and its gradient are dropped, while only a small penalty on the second gradient of the deformation is added to stabilize the simulations. Thus, all of the numerical simulations should be driven towards the analytical solutions. That is exactly what we observe.
Figs. 7, 8(b) and (f) plot the effective deformations obtained from the simulations. Notice that the responses are genuinely large deformation in each case, with the characteristic out-of-plane displacements on the order of $\sim 0.2$-$0.5$ for domains of characteristic length $\sim 1.4$-$1.8$. Yet the errors in the deformation are negligible by comparison. Figs. 7, 8(c) and (g) plot normed differences between the analytical and simulated effective deformations. The errors are on the order of $0.1$-$1\%$, depending on the simulation. The twist case performs especially well in this comparison. Figs. 7, 8(d) and (h) also compare the angle fields. For reference, this field varies in magnitude from 00 to $0.5$ in the typical simulation. From the figures, we see that its errors are on the order of $5$-$20$% in places, especially towards the boundaries of the pattern in the Miura cases. However, in most of the domain, the errors are far below $1\%$. The deformations are perhaps better approximated because of our choice of boundary conditions for the simulations. In any case, the simulations and analytical solutions are mostly in excellent agreement.
5.3 Pinching deformations
Having validated the numerical method in the previous two sections, we now highlight its versatility by exploring inhomogeneous boundary conditions for which analytical solutions are difficult to come by.
[…]
Plots of the normalized force versus the pinching strain are shown in Fig. 9(a) for the Miura and Fig. 10(a) for the Eggbox. The force is calculated numerically using a variational method detailed in Appendix E; it is scaled by the folding modulus $d_{3}$ to give the normalized force in the plots. We also reproduce from Section 5.1 the pure mechanism results on these plots, as well as the case where the boundary condition $\mathbf{y}_{h}=\mathbf{A}(\overline{\theta})\mathbf{x}$ is applied to the full left and right parts of the boundary of $\partial\Omega$. Figs. 9 and 10(b) display the variation of the angle field versus the pinching strain in each simulation and for each boundary condition, while Figs. 9 and 10(c-e) highlight representative examples of the pinching simulations at large strains. Finally, Table 1 compares the different energy contributions for the simulations at $20\%$ strain intervals.
Miura Pinching Simulations 20% 40 % 60% 80%
[…]
Eggbox Pinching Simulations 20% 40 % 60% 80%
[…]
Table 1: Comparison of the different energy contributions during the pinching simulations for $c_{1}=5,d_{1}=d_{2}=10^{-2},d_{3}=0.1$ and $L_{\Omega}=1.7$ for Miura and $=\sqrt{2}$ for Eggbox. Snapshots of each energy are taken at $20\%$ strain intervals.
Figure 9: Pinching simulations for Miura origami. (a) The normalized force versus compressive strain for the pinching simulations (black) compared to that of the pure mechanism (red) and the case where the full left and right boundaries are displaced. (b) The angle envelop (minimum to maximum) at each value of compressive strain. The angle is constant for the pure mechanism. (c) Reference configuration prior to pinching; (d) 40% compressive strain; (e) 80% compressive strain.
As the boundaries are pinched, the folds must actuate in regions around $\Gamma_{\text{d}}=\Gamma_{L}\cup\Gamma_{R}$ to accommodate the underlying compression. This actuation, being energetically costly, is then gradually relaxed in a manner regularized by the $\nabla\theta_{h}$ and $\nabla\nabla\mathbf{y}_{h}$ terms in the energy, yielding the bowtie shapes shown. Evidently, all of this is done without incurring too much bulk energy: notably, the bulk energy modulus $c_{1}$ is $50$ times larger that the folding modulus $d_{3}$, yet the bulk energy contribution is consistently one order of magnitude lower than the dominant folding energy (see Table 1). Consequently, much like the pure mechanism, the folding modulus $d_{3}$ sets the force for the pinching simulations. Intriguingly, the total force to pinch the sample at a given engineering strain is comparable to the pure mechanism, and in fact larger for most strains. This is likely due to the basic nature of the boundary forces in our generalized continuum. They consist of two terms — a traditional Piola-Kirchhoff stress $\mathbf{P}(\mathbf{y}_{h},\theta_{h})$ and a generalized stress proportional to $\nabla\mathcal{H}(\mathbf{y}_{h},\theta_{h})$ (see Appendix E). The former is comparable for both cases; the stress is slightly higher for the pinching simulations, but covers only $20\%$ of the boundary. The latter, however, is zero for the pure mechanism since the deformation and actuation in this case is uniform. It is alternatively quite high in the pinching simulations since the higher-order gradients contributing to this term are magnified near the compressed boundaries.
Figure 10: Pinching simulations for Eggbox origami. (a) The normalized force versus compressive strain for the pinching simulations (black) compared to that of the pure mechanism (red) and the case where the full left and right boundaries are displaced. (b) The angle envelop (minimum to maximum) at each value of compressive strain. The angle is constant for the pure mechanism. (c) Reference configuration prior to pinching; (d) 40% compressive strain; (e) 80% compressive strain.
Beyond these general observations, which hold similarly for the Miura and Eggbox cases, there are distinct differences in the two pinching simulations. Notice that the angle variation at each value of compressive strain is much higher for the Miura case (Fig. 9(b)) compared to the Eggbox (Fig. 10(b)). The contours illustrating this variation are also qualitatively quite different. For the Miura in Fig. 9(c-e), they form ellipses around the compressive loading on $\Gamma_{L}$ and $\Gamma_{R}$. In Eggbox, there is a distinct diamond shape of nearly constant actuation in the bulk of the sample in Fig. 10(c-e). Accordingly, the shape of the Miura is a pronounced bowtie, while that of the Eggbox more closely resembles the pure mechanism.
[…]
$\mathbf{I}(\mathbf{y}_{\text{g}})=\mathbf{A}^{T}(\theta_{\text{g}})\mathbf{A}(\theta_{\text{g}})=\begin{pmatrix}\tfrac{\lambda^{2}_{1}(\theta_{\text{g}})}{\lambda^{2}_{1}(0)}&0\\ 0&\tfrac{\lambda^{2}_{2}(\theta_{\text{g}})}{\lambda^{2}_{2}(0)}\end{pmatrix}\quad\text{ and }\quad\mathbf{II}(\mathbf{y}_{\text{g}})=\mathbf{0},$ (66)
since the bulk energy in each simulation is negligible compared to its modulus. An application of Gauss’s theorem egregium to these formulas then furnishes a PDE governing $\theta_{\text{g}}$ of the form
$\partial_{1}\Big{(}\frac{\lambda_{1}(0)\lambda_{2}^{\prime}(\theta_{\text{g}})}{\lambda_{2}(0)\lambda_{1}(\theta_{\text{g}})}\partial_{1}\theta_{\text{g}}\Big{)}+\partial_{2}\Big{(}\frac{\lambda_{2}(0)\lambda_{1}^{\prime}(\theta_{\text{g}})}{\lambda_{1}(0)\lambda_{2}(\theta_{\text{g}})}\partial_{2}\theta_{\text{g}}\Big{)}=0.$ (67)
This second-order PDE is elliptic if $\lambda_{1}^{\prime}(\theta)\lambda_{2}^{\prime}(\theta)>0$ for all $\theta$ and hyperbolic if $\lambda_{1}^{\prime}(\theta)\lambda_{2}^{\prime}(\theta)<0$ for all $\theta$. In other words, it is elliptic for patterns like Miura origami that are auxetic and hyperbolic for those like Eggbox that are not auxetic.
Ellipticity versus hyperbolicity is a standard classification in PDE theory, with clear and general ramifications. Elliptic PDEs describe a decay in the field variables away from loads, highlighted by the fact that the maximum and minimum of the field variables always occurs on the boundary and never on the interior unless they are constant. [^11] This is called the strong maximum principle of Elliptic PDEs. The Miura in Fig. 9(c-e) perfectly reflects these properties — the angle field $\theta_{h}$ is largest at the loaded boundary, decays away along elliptic contours from this boundary, and is at its minimum at the four corners. Hyperbolic PDEs are the diametric opposite. They describe wave-like phenomena, where the field variable instead propagates from the loaded boundary to the interior of the domain along characteristic curves. Characteristics also cannot intersect, making the solutions to hyperbolic PDEs generally more rigid to Dirichlet boundary data. The Eggbox in 10(c-e) reflects these properties — a large diamond region of nearly constant actuation on the interior of the domain recedes to the corners through “fan-like” characteristics; the pattern also displays far less heterogeneity, thus less “flexibility” in a sense, than its Miura counterpart.
An astute observer may notice that the Eggbox case does not completely fit the description of a hyperbolic PDE without a little embellishment — the actuation is still largest at the loaded boundaries and smallest at the corners, like in the Miura case. To explain, we must remember that $\theta_{h}$ only approximates $\theta_{\text{g}}$. In fact, there is a universal elliptic part to our generalized continuum model. The governing equations in (40) include the equilibrium condition $\nabla\cdot\mathbf{j}(\theta_{h})=q(\mathbf{y}_{h},\theta_{h})$, which is elliptic in $\theta_{h}$ since $\nabla\cdot\mathbf{j}(\theta_{h})\sim\nabla^{2}\theta_{h}$. This complicates the qualitative behavior of the Eggbox. In particular, $\theta_{h}$ must approximate an angle field $\theta_{\text{g}}$ associated to the hyperbolic equation in (67) while simultaneously solving an elliptic equilibrium equation.
5.4 Bending deformations under transverse loads
For our final set of examples, we examine the behavior of Miura and Eggbox origami under transverse loads. The basic setup is shown in Fig. 11(a). We apply a uniform transverse load $\overline{\mathbf{b}}=b\mathbf{e}_{3}$ (for a constant $b$) to the effective domain $\Omega$ of each pattern and pin the boundaries via the Dirichlet conditions $\mathbf{y}_{h}=\mathbf{x}$ on $\Gamma_{L}$ and $\Gamma_{R}$, which are centered on the sample. We study this boundary value problem in the cases that $\Gamma_{L,R}$ take up $5\%$, $10\%$, and $20\%$ of the left and right boundaries, respectively. By partially pinning these boundaries, as opposed to say fully pinning them, each pattern can accommodate the load via doubly-curved shapes with minimal geometric frustration. The choice of parameters in these simulations is the same as the pinching case: $c_{1}=5,c_{2}=1,d_{1}=d_{2}=10^{-2}$, $d_{3}=0.1$ and $L_{\Omega}=1.7$ for Miura and $=\sqrt{2}$ for Eggbox.
Fig. 11(b) plots the average out-of-plane displacement $\bar{w}=\frac{1}{|\Omega|}\int_{\Omega}\mathbf{y}\cdot\mathbf{e}_{3}\mathop{}\!\mathrm{d}x$ versus the load $b$ for each simulation of each pattern under each boundary condition. Fig. 11(c) shows three representative deformed configurations for Miura in the $10\%$ pinned case; Fig. 11(d) shows analogous such shapes for the Eggbox, again in the $10\%$ case.
Figure 11: Bending simulations under transverse load. (a) The pattern is pinned at $\Gamma_{L}$ and $\Gamma_{R}$ and subject to the uniform load $b$. (b) Plot of the average displacement $\bar{w}=\frac{1}{|\Omega|}\int_{\Omega}\mathbf{y}\cdot\mathbf{e}_{3}\mathop{}\!\mathrm{d}x$ versus the load. (c) Deformation profiles for Miura and (d) for Eggbox at representative points on the loading curves, labeled in (b). Each of these profiles is for the case where $10\%$ of the left and right boundaries is pinned.
At the early stages of loading, both Miura and Eggbox deform quite softly, with the stored energy is dominated by the bending term $\int_{\Omega}d_{1}|\nabla\nabla\mathbf{y}|^{2}\mathop{}\!\mathrm{d}x$. However, geometric frustration sets in as the loading progresses. At the tail end of the simulations, the two bulk energy terms $\int_{\Omega}W_{1}(\theta,\mathbf{II}(\mathbf{y}))\mathop{}\!\mathrm{d}x$ and $\int_{\Omega}W_{2}(\theta,\mathbf{II}(\mathbf{y}))\mathop{}\!\mathrm{d}x$ become comparable to the bending term. This transition — from negligible to prominent bulk stored energy — coincides with the universal stiffening response shown in the plots, and is likely its cause. Another notable, if unsurprising, feature of the force plot is that the patterns are stiffer for the $20\%$ pinned boundary compared to the $10\%$ and to the $5\%$ cases. This is easily explained by noticing that each pattern needs to flatten out in the $x_{2}$-direction toward the $\Gamma_{L,R}$ boundaries, and thus become approximately singly curved in small subregions of $\Omega$ that contain these boundaries. A significant bulk energy contribution $W_{2}(\theta,\mathbf{II}(\mathbf{y}))$ necessarily arises in these regions because $W_{2}$ biases the patterns to be doubly-curved. The last and most surprising observation from these plots is that the bending stiffness of Eggbox is much larger than that of Miura. This is true even though the in-plane stiffness of the Miura greatly exceeds the Eggbox for both the pure mechanism and the pinching simulations, illustrated previously.
We now spend some time analyzing and discussing the differences in bending stiffness between the two examples. As a basic heuristic, we believe this difference is essentially explained by the fact that the saddle-like shapes preferred by the Miura can do more work $\int_{\Omega}b\mathbf{y}\cdot\mathbf{e}_{3}\mathop{}\!\mathrm{d}x$ at each value of $b$ than the analogous shapes for the Eggbox. Notice that the the cap-like shapes common to Eggbox are not observed in the simulations. Instead, the pattern forms a cap in its center but then reverses its curvature towards the boundary to induce more vertical displacement and thus more work. This geometric frustration underlies the Eggbox’s much stiffer response.
We can quantitatively investigate this heuristic, at least in a linearized regime $|b|\ll 1$ where the analysis is more tractable. The idea is to minimize the total potential energy at leading order for bending-type deformations of the form
$\mathbf{y}_{b}(\mathbf{x})=\begin{pmatrix}\mathbf{x}\\ bw(\mathbf{x})\end{pmatrix}+O(b^{2}),\quad\theta_{b}(\mathbf{x})=O(b^{2}),$ (68)
with $b$ playing the role of the small parameter in the asymptotics. Notice that the out-of-plane displacement $bw$ is taken to be much larger than the planar components, as one would expect under small transverse loading. [^12] At first glance, it’s not obvious that the actuation should scale as $b^{2}$ for small $b$. However, after expanding $\theta_{b}$ in powers of $b$ via $\theta_{b}=\delta\theta_{0}+b\delta\theta_{1}+b^{2}\delta\theta_{2}+\ldots$, one will find upon energy minimization that $\delta\theta_{0}=\delta\theta_{1}=0$. Substituting these deformations into the total energy leads to
$E(\mathbf{y}_{b},\theta_{b})=b^{2}\Big{(}c_{2}L_{\Omega}^{2}\int_{\Omega}\Big{(}\frac{\lambda_{1}^{\prime}(0)}{\lambda_{1}(0)}\partial_{2}\partial_{2}w+\frac{\lambda_{2}^{\prime}(0)}{\lambda_{2}(0)}\partial_{1}\partial_{1}w\Big{)}^{2}\mathop{}\!\mathrm{d}x+d_{1}L_{\Omega}^{2}\int_{\Omega}|\nabla\nabla w|^{2}\mathop{}\!\mathrm{d}x-\int_{\Omega}w\mathop{}\!\mathrm{d}x\Big{)}+o(b^{2}).$ (69)
As the moduli satisfy $c_{2}\gg d_{1}$ in the typical case, it is reasonable to impose the PDE constraint
$\frac{\lambda_{1}^{\prime}(0)}{\lambda_{1}(0)}\partial_{2}\partial_{2}w+\frac{\lambda_{2}^{\prime}(0)}{\lambda_{2}(0)}\partial_{1}\partial_{1}w=0$ (70)
upon minimization. This requirement results in a small correction to the overall minimum energy that can be neglected for our purposes. It follows that
[…]
for all sufficiently small $b$. Note that $b$ is the load per unit area and $\frac{1}{|\Omega|}\int_{\Omega}bw\mathop{}\!\mathrm{d}x$ is the average displacement under this load. Thus, if $w^{\star}$ is a minimizer to the latter, then the normalized bending stiffness relevant to the plots in Fig. 11(b) is the ratio
$K_{\text{bend}}=\frac{1}{\frac{1}{|\Omega|}\int_{\Omega}w^{\star}\mathop{}\!\mathrm{d}x}.$ (72)
Instead of trying to perform the PDE-constrained optimization in (71) — which is challenging and likely overkill for the task of explaining the differences in stiffness between the two examples — we pursue a minimization within an ansatz. First, to simplify the analysis, we replace the reference configuration in (52) with one centered at the origin $\Omega=(-\lambda_{1}(0)/2,\lambda_{1}(0)/2)\times(-\lambda_{2}(0)/2,\lambda_{2}(0)/2)$ so that the minimizer $w^{\star}$ to (71) is even in $x_{1}$ and $x_{2}$. The most general fourth order-polynomial that has this symmetry and solves the PDE in (70) is
$w_{\text{poly}}(x_{1},x_{2})=w_{0}+\kappa_{0}\Big{(}\frac{\lambda_{1}^{\prime}(0)}{\lambda_{1}(0)}x_{1}^{2}-\frac{\lambda_{2}^{\prime}(0)}{\lambda_{2}(0)}x_{2}^{2}\Big{)}+\tau_{0}\Big{(}\frac{\lambda^{\prime}_{1}(0)\lambda_{2}(0)}{6\lambda_{2}^{\prime}(0)\lambda_{1}(0)}x_{1}^{4}-x_{1}^{2}x_{2}^{2}+\frac{\lambda^{\prime}_{2}(0)\lambda_{1}(0)}{6\lambda_{1}^{\prime}(0)\lambda_{2}(0)}x_{2}^{4}\Big{)}$ (73)
for constants $w_{0},\kappa_{0}$ and $\tau_{0}$. We supply the constant $w_{0}=w_{0}(\kappa_{0},\tau_{0})$ in this formula by imposing the condition $w_{\text{poly}}(\lambda_{1}(0)/2,0)=0$, which is a proxy for the boundary data $w=0$ on $\Gamma_{L,R}$ in the minimization in (71). Hence, a “back-of-the-envelope” prediction of the normalized bending stiffness is
$K_{\text{bend}}\approx\frac{1}{\frac{1}{|\Omega|}\int_{\Omega}w_{\text{poly}}^{\star}\mathop{}\!\mathrm{d}x},$ (74)
where $w_{\text{poly}}^{\star}$ solves the simple minimization problem
$w^{\star}_{\text{poly}}=\operatorname*{argmin}_{\kappa_{0},\tau_{0}}\int_{\Omega}\Big{(}d_{1}L_{\Omega}^{2}|\nabla\nabla w_{\text{poly}}|^{2}-w_{\text{poly}}\Big{)}\mathop{}\!\mathrm{d}x.$ (75)
As is typical with these calculations, the formulas for the minimizing coefficients and the approximate stiffness are lengthy expressions of $\lambda_{1,2}(0)$ and $\lambda_{1,2}^{\prime}(0)$, which we do not report. After substituting for their explicit values using (50) and (51), the stiffness is
$\frac{1}{\frac{1}{|\Omega|}\int_{\Omega}w_{\text{poly}}^{\star}\mathop{}\!\mathrm{d}x}=d_{1}L_{\Omega}^{2}\begin{cases}31.9&\text{ for Miura}\\ 322.2&\text{ for Eggbox}.\end{cases}$ (76)
Then, using the values of $d_{1}$ and $L_{\Omega}$ from the simulations, we calculate from (76) a normalized stiffness of $0.92$ for Miura and $6.44$ for Eggbox. The corresponding simulated values at the origin are $0.72$ and $4.44$ for Miura and Eggbox, respectively. So this expression gets the trends right, namely that Eggbox is stiffer than the Miura under small transverse loads, but also provides reasonable quantitative agreement in spite of all the simplifications.
One nice feature of capturing the trends in a simplified setting is that it enables design optimization without taxing a supercomputer. As a quick example, we can replace the expressions for $\lambda_{i}(\theta)$ in (50) and (51) with $\lambda_{i}(\theta+\psi)$ for an angle field $\psi$, which alters the reference folded state of the unit cells, and then optimize the stiffness as a function of $\psi$ using our simplified model. In the Miura case, the stiffness is unsurprisingly monotonic — it achieves its minimum at the fully flat state and becomes infinite at the fully folded state. The Eggbox too has a stiffness that approaches infinity at the fully folded state. However, it is most compliant at the nontrivial value $\psi\approx 0.21$. Thus, sharply corrugated patterns have high stiffness, while extremal compliance is a more delicate matter, especially in non-Euclidean cases like the Eggbox.
We end by pointing out an intriguing connection between pinching and linearized bending in these metamaterials: Both situations are well characterized by underlying second-order PDEs whose “PDE type”, and thus its basic qualitative properties, is completely dictated by the pattern’s auxeticity. Specifically, the actuation field in the pinching simulations is well approximated by the PDE for $\theta_{\text{g}}$ in (67), while the out-of-plane displacement $pw$ under small transversed load is constrained by (70). Both PDEs are elliptic for Miura Origami (or really any auxetic parallelogram origami) and hyperbolic for Eggbox (any pattern that is not auxetic). More generally, they emerge as different analytical case-studies of the geometric constraints $\mathbf{I}(\mathbf{y})=\mathbf{A}^{T}(\theta)\mathbf{A}(\theta)$ and $\big{(}\mathbf{u}(\theta)\cdot\mathbf{u}^{\prime}(\theta)\big{)}\tilde{\mathbf{v}}_{0}\cdot\mathbf{II}(\mathbf{y})\tilde{\mathbf{v}}_{0}=-\big{(}\mathbf{v}(\theta)\cdot\mathbf{v}^{\prime}(\theta)\big{)}\tilde{\mathbf{u}}_{0}\cdot\mathbf{II}(\mathbf{y})\tilde{\mathbf{u}}_{0}$ that underly our modeling framework. In fact, we anticipate that any such case-study of these constraints will lead to PDE compatibility conditions that illuminate an “elliptic $=$ auxetic” versus “hyperbolic $=$ non-auxetic” dichotomy in the response, although we leave this point, to be examined more carefully, for another time.
6 Conclusions
This paper brought together rigorous theory, modeling, numerical method development, simulations, and mechanical analysis to address the challenging problem of understanding the effective elastic behavior of parallelogram origami. We first described a pair of coarse-graining rules that produced effective geometric constraints for origami soft modes and a corresponding plate energy for their elasticity. Then, we introduced a simplified elastic model that drove the pattern’s effective deformation and actuation towards the geometric constraints of this theory, while accounting for higher-order sources of elasticity — the plate energy — through an appropriate regularization. Next, we provided a finite element formulation of this model using the $C^{0}$ interior penalty method. Finally, through numerical implementation in Firedrake, we demonstrated a model and numerical framework capable of simulating the effective behavior of parallelogram origami metamaterials across a variety of loading conditions
Challenges and opportunities remain. It would certainly be appealing to compare our model and numerics to experiments, both qualitatively and quantitatively. We anticipate that, with proper tuning of the parameters, our approach will capture the gross-shape change observed in experiments across designs and loading condition. However, matching the loading curves may be more challenging. Potential improvements are readily available. We took a purposefully simple approach to modeling the higher-order effects in the theory. It is possible, for instance, to replace the $\int_{\Omega}(d_{1}L_{\Omega}^{2}|\nabla\nabla\mathbf{y}|^{2}+d_{2}L_{\Omega}^{2}|\nabla\theta|^{2})\mathop{}\!\mathrm{d}x$ term in the energy with the more complicated but asymptotically correct effective plate energy in (17). A mathematical challenge is to prove the existence of minimizers in this setting — the proof in Theorem 1 makes key use of the fact that the $L^{2}$-norm of $\nabla\nabla\mathbf{y}$ is controlled by the elastic energy, and will need to be adapted. Another natural modification is to replace the actuation energy $\int_{\Omega}d_{3}|\theta|^{2}\mathop{}\!\mathrm{d}x$ with a weighted average $\int_{\Omega}d_{3}\sum_{i=1,\ldots,4}\lambda_{i}|\gamma_{i}(\theta)|^{2}\mathop{}\!\mathrm{d}x$, where $\gamma_{i}\colon(\theta^{-},\theta^{+})\rightarrow\mathbb{R}$ parameterize the change in the four distinct dihedral angles of the unit cell and $\lambda_{i}$ are the weights. [^13] Presumably, each $\lambda_{i}$ should be determined geometrically by the corresponding creases perimeter or some notion of its influence area. Whether such modifications are “worth the bang for your buck” will require synergistic modeling and experimental efforts going forward.
It would also be appealing to use our framework to explore large swaths design space of parallelogram origami, beyond the canonical Miura and Eggbox origami. Importantly, the basic setup of our model allows for any parallelogram origami design. So all that needs to be done to begin this exploration is to feed the appropriate design dependent parameterizations of $\mathbf{u}(\theta)$ and $\mathbf{v}(\theta)$ into the model. Another intriguing research direction is to study “locally periodic” origami patterns, which are obtained by slowly varying the design parameters of the origami cells and often show an uncanny ability to approximate a rich variety of surfaces on folding [14, 18, 53]. The lattice vectors in our model $\mathbf{u}_{0}\equiv\mathbf{u}_{0}(\mathbf{d}),\ldots,\mathbf{v}(\theta)\equiv\mathbf{v}(\theta,\mathbf{d})$ implicitly depend on design angles and lengths, labeled $\mathbf{d}$. In a locally periodic setting, $\mathbf{d}=\mathbf{d}(\mathbf{x})$ spatially varies in the reference domain $\Omega$, introducing a material anisotropy to the model. We also anticipate the emergence of an effective elasticity proportional to $\nabla\mathbf{d}$ on coarse-graining the mechanical response of these patterns, although pinning down the exact details of this elasticity seems far from trivial.
As a final point, flexible mechanical metamaterials are appealing in applications both for their large design space and their highly design dependent nonlinear mechanical response. It is thus important to have modeling tools that efficiently map the relationship between design and bulk mechanical properties in these systems. Our results here, and those of our prior works [64, 67, 68], show that homogenization is a predictive and useful tool in this setting, while also pushing the boundaries of the theory of elasticity in new and exciting directions. We hope our work will serve as an exemplar for further exploration along these lines.
Data availability
The code for all numerical simulations in this paper is available on GitHub at:
https://github.com/HuUSC/ContinuumModelOri.git
Acknowledgments
H.X. and P.P. acknowledge support from the Army Research Office (ARO-W911NF2310137). P.P. also acknowledges support from the National Science Foundation (CMMI-CAREER-2237243).
F.M. acknowledges support from the National Science Foundation (DMS-2409926).
Appendix A Proof of the existence of minimizers for the constitutive model
This section proves Theorem 1 on the existence of minimizers to the energy in Section 3. We break the proof up into several steps.
Step 1. Extensions of the energy densities. We find it useful to express the internal energy $E_{\text{int}}(\mathbf{y},\theta)$ defined by (19-24) as
$E_{\text{int}}(\mathbf{y},\theta)=\int_{\Omega}W^{\varepsilon}(\theta,\nabla\theta,\nabla\mathbf{y},\nabla\nabla\mathbf{y})\mathop{}\!\mathrm{d}x,\quad(\mathbf{y},\theta)\in V^{\varepsilon}$ (77)
for an energy density $W^{\varepsilon}\colon\mathbb{R}\times\mathbb{R}^{2}\times\mathbb{R}^{3\times 2}\times\mathbb{R}^{3\times 2\times 2}\rightarrow\mathbb{R}$ that is extended to all of space and continuous in every argument. In particular, we construct a $W^{\varepsilon}$ that coincides with the original energy density for $E_{\text{int}}(\mathbf{y},\theta)$ whenever $(\mathbf{y},\theta)\in V^{\varepsilon}$, but deviates from this density in a continuous way otherwise. This modification allows us to appeal to general results in the direct method of the calculus of variations to establish the existence of minimizers of the energy. The construction makes use of the $\varepsilon$-dependence of the set $V^{\varepsilon}$.
To begin, recall from Section 2.1 that $\mathbf{u}(\theta),\mathbf{v}(\theta),\mathbf{A}(\theta),\mathbf{u}^{\prime}(\theta),\mathbf{v}^{\prime}(\theta)$ are well-defined and smooth on the interval $(\theta^{-},\theta^{+})$. However, $\mathbf{u}^{\prime}(\theta)$ and/or $\mathbf{v}^{\prime}(\theta)$ can blow up as $\theta\rightarrow\theta^{+}$ and/or $\theta^{-}$ for some parallelogram origami designs. As we are really focused on the behavior of this actuation on the interval $[\theta^{-}+\varepsilon,\theta^{+}-\varepsilon]$ per $V^{\varepsilon}$, we introduce convenient $\varepsilon$-extensions of these functions $\mathbf{u}_{\varepsilon}(\theta)$, $\mathbf{v}_{\varepsilon}(\theta)$, $\mathbf{A}_{\varepsilon}(\theta)$, $[\mathbf{u}^{\prime}]_{\varepsilon}(\theta)$ and $[\mathbf{v}^{\prime}]_{\varepsilon}(\theta)$ as follows. Define $\mathbf{u}_{\varepsilon}\colon\mathbb{R}\rightarrow\mathbb{R}^{3}$ as
$\mathbf{u}_{\varepsilon}(\theta):=\begin{cases}\mathbf{u}(\theta)&\text{ if }\theta\in[\theta^{-}+\varepsilon,\theta^{+}-\varepsilon]\\ \mathbf{u}(\theta^{-}+\varepsilon)&\text{ if }\theta<\theta^{-}+\varepsilon\\ \mathbf{u}(\theta^{+}-\varepsilon)&\text{ if }\theta>\theta^{+}-\varepsilon,\end{cases}$ (78)
and $\mathbf{v}_{\varepsilon}\colon\mathbb{R}\rightarrow\mathbb{R}^{3}$ analogously. Then, define $\mathbf{A}_{\varepsilon}\colon\mathbb{R}\rightarrow\mathbb{R}^{3\times 2}$ as the unique linear transformation that satisfies
$\mathbf{A}_{\varepsilon}(\theta)\tilde{\mathbf{u}}_{0}=\mathbf{u}_{\varepsilon}(\theta)\quad\text{ and }\quad\mathbf{A}_{\varepsilon}(\theta)\tilde{\mathbf{v}}_{0}=\mathbf{v}_{\varepsilon}(\theta),$ (79)
and note that this definition implies that $\mathbf{A}_{\varepsilon}(\theta)=\mathbf{A}(\theta)$ for all $\theta\in[\theta^{-}+\varepsilon,\theta^{+}-\varepsilon]$. Finally, define $[\mathbf{u}^{\prime}]_{\varepsilon}\colon\mathbb{R}\rightarrow\mathbb{R}$ similar to (78) via
$[\mathbf{u}^{\prime}]_{\varepsilon}(\theta):=\begin{cases}\mathbf{u}^{\prime}(\theta)&\text{ if }\theta\in[\theta^{-}+\varepsilon,\theta^{+}-\varepsilon]\\ \mathbf{u}^{\prime}(\theta^{-}+\varepsilon)&\text{ if }\theta<\theta^{-}+\varepsilon\\ \mathbf{u}^{\prime}(\theta^{+}-\varepsilon)&\text{ if }\theta>\theta^{+}-\varepsilon,\end{cases}$ (80)
and define $[\mathbf{v}^{\prime}]_{\varepsilon}\colon\mathbb{R}\rightarrow\mathbb{R}^{3}$ analogously. Note that these extensions are not chosen to be $\mathbf{u}_{\varepsilon}^{\prime}(\theta),\mathbf{v}^{\prime}_{\varepsilon}(\theta)$ because the latter are discontinuous at $\theta=\theta^{\mp}\pm\varepsilon$.
Next, we address the fact that the surface normal $\mathbf{n}(\mathbf{y})$ in (12) is not well-defined when $\partial_{1}\mathbf{y}\times\partial_{2}\mathbf{y}=\mathbf{0}$. Embracing again the $\varepsilon$-dependence in $V^{\varepsilon}$, we regularize the surface normal via the continuous vector field $\bm{\nu}_{\varepsilon}\colon\mathbb{R}^{3\times 2}\rightarrow\mathbb{R}^{3}$ defined by
$\bm{\nu}_{\varepsilon}(\mathbf{F}):=\begin{cases}\frac{\mathbf{F}\mathbf{e}_{1}\times\mathbf{F}\mathbf{e}_{2}}{|\mathbf{F}\mathbf{e}_{1}\times\mathbf{F}\mathbf{e}_{2}|}&\text{ if }|\mathbf{F}\mathbf{e}_{1}\times\mathbf{F}\mathbf{e}_{2}|\geq\varepsilon\\ \varepsilon^{-1}\mathbf{F}\mathbf{e}_{1}\times\mathbf{F}\mathbf{e}_{2}&\text{ otherwise},\end{cases}$ (81)
where $\{\mathbf{e}_{1},\mathbf{e}_{2}\}$ denotes the standard basis on $\mathbb{R}^{2}$. It follows that $\bm{\nu}_{\varepsilon}(\nabla\mathbf{y})=\mathbf{n}(\mathbf{y})$ whenever $|\partial_{1}\mathbf{y}\times\partial_{2}\mathbf{y}|\geq\varepsilon$. Likewise, we address the potential blowup of $1/\sqrt{\det\mathbf{I}(\mathbf{y})}$ in $W_{1}$ of (LABEL:eq:W12Def) as follows. Noting that $\det\mathbf{I}(\mathbf{y})=|\partial_{1}\mathbf{y}\times\partial_{2}\mathbf{y}|^{2}$, we define
$j_{\varepsilon}(\mathbf{F})=\begin{cases}|\mathbf{F}\mathbf{e}_{1}\times\mathbf{F}\mathbf{e}_{1}|&\text{ if }|\mathbf{F}\mathbf{e}_{1}\times\mathbf{F}\mathbf{e}_{2}|\geq\varepsilon\\ \varepsilon&\text{ otherwise}.\end{cases}$ (82)
It follows that $1/j_{\varepsilon}(\mathbf{F})$ is continuous and $\sqrt{\det\mathbf{I}(\mathbf{y})}=j_{\varepsilon}(\nabla\mathbf{y})$ whenever $|\partial_{1}\mathbf{y}\times\partial_{2}\mathbf{y}|\geq\varepsilon.$
We now define $W^{\varepsilon}$. For this purpose and throughout the remainder of this section, the variables $\mathbf{g}\in\mathbb{R}^{2},\mathbf{F}\in\mathbb{R}^{3\times 2},\mathcal{F}\in\mathbb{R}^{3\times 2\times 2},\mathcal{F}_{\alpha\beta}\in\mathbb{R}^{3}$ denote placeholders for $\nabla\theta,\nabla\mathbf{y},\nabla\nabla\mathbf{y}$, $\partial_{\alpha}\partial_{\beta}\mathbf{y}$, respectively. Hence, we take $W^{\varepsilon}\colon\mathbb{R}\times\mathbb{R}^{2}\times\mathbb{R}^{3\times 2}\times\mathbb{R}^{3\times 2\times 2}\rightarrow\mathbb{R}$ to satisfy
$W^{\varepsilon}(\theta,\mathbf{g},\mathbf{F},\mathcal{F}):=W_{1}^{\varepsilon}(\theta,\mathbf{F})+W_{2}^{\varepsilon}(\theta,\mathbf{F},\mathcal{F})+d_{1}|\mathcal{F}|^{2}+d_{2}|\mathbf{g}|^{2}+d_{3}|\theta|^{2}$ (83)
for energy densities $W_{1,2}^{\varepsilon}$ that are continuous extensions of $W_{1,2}$ in (LABEL:eq:W12Def) and defined by
$W_{1}^{\varepsilon}(\theta,\mathbf{F}):=\frac{c_{1}}{j_{\varepsilon}(\mathbf{F})}\big{|}\mathbf{F}^{T}\mathbf{F}-\mathbf{A}_{\varepsilon}^{T}(\theta)\mathbf{A}_{\varepsilon}(\theta)\big{|}^{2},$ (84)
$W_{2}^{\varepsilon}(\theta,\mathbf{F},\mathcal{F}):=\frac{c_{2}L_{\Omega}^{2}}{|\tilde{\mathbf{u}}_{0}|^{4}|\tilde{\mathbf{v}}_{0}|^{4}}\Big{(}\big{[}\mathbf{v}_{\varepsilon}(\theta)\cdot[\mathbf{v}^{\prime}]_{\varepsilon}(\theta)\big{]}\big{[}\tilde{\mathbf{u}}_{0}\cdot\mathbf{K}_{\varepsilon}(\mathbf{F},\mathcal{F})\tilde{\mathbf{u}}_{0}\big{]}+\big{[}\mathbf{u}_{\varepsilon}(\theta)\cdot[\mathbf{u}^{\prime}]_{\varepsilon}(\theta)\big{]}\big{[}\tilde{\mathbf{v}}_{0}\cdot\mathbf{K}_{\varepsilon}(\mathbf{F},\mathcal{F})\tilde{\mathbf{v}}_{0}\big{]}\Big{)}^{2},$
where $\mathbf{K}_{\varepsilon}\colon\mathbb{R}^{3\times 2}\times\mathbb{R}^{3\times 2\times 2}\rightarrow\mathbb{R}^{2\times 2}_{\text{sym}}$ is the tensor field
$\mathbf{K}_{\varepsilon}(\mathbf{F},\mathcal{F}):=\begin{pmatrix}\mathcal{F}_{11}\cdot\bm{\nu}_{\varepsilon}(\mathbf{F})&\mathcal{F}_{12}\cdot\bm{\nu}_{\varepsilon}(\mathbf{F})\\ \mathcal{F}_{12}\cdot\bm{\nu}_{\varepsilon}(\mathbf{F})&\mathcal{F}_{22}\cdot\bm{\nu}_{\varepsilon}(\mathbf{F})\end{pmatrix}.$ (85)
As with the surface normal, $\mathbf{K}_{\varepsilon}(\nabla\mathbf{y},\nabla\nabla\mathbf{y})=\mathbf{II}(\mathbf{y})$ whenever $|\partial_{1}\mathbf{y}\times\partial_{2}\mathbf{y}|\geq\varepsilon$; it is also linear in $\mathcal{F}$. This completes our construction of the desired extended and continuous energy density $W^{\varepsilon}$ in (77).
Step 2. Coercivity, growth, and convexity. We turn to establishing several pointwise properties of $W^{\varepsilon}$, including upper and lower bounds and convexity in certain arguments. These properties are key to applying the direct method of the calculus of variations to prove the existence of minimizers.
Our first result concerns estimates on the energy density $W_{1}^{\varepsilon}$ in (LABEL:eq:W12Eps).
Lemma 1.
There are constants $0<c_{1,\varepsilon}<C_{1,\varepsilon}$ such that
$c_{1,\varepsilon}|\mathbf{F}|^{2}-\frac{1}{c_{1,\varepsilon}}\leq W^{\varepsilon}_{1}(\theta,\mathbf{F})\leq C_{1,\varepsilon}(|\mathbf{F}|^{4}+1)\quad\text{ for all }\quad(\theta,\mathbf{F})\in\mathbb{R}\times\mathbb{R}^{3\times 2}.$ (86)
Proof.
For the upper bound, observe by standard estimates that
$W^{\varepsilon}_{1}(\theta,\mathbf{F})=\frac{c_{1}}{j_{\varepsilon}(\mathbf{F})}|\mathbf{F}^{T}\mathbf{F}+\mathbf{A}_{\varepsilon}^{T}(\theta)\mathbf{A}_{\varepsilon}(\theta)|^{2}\leq 2c_{1}\varepsilon^{-1}\big{(}|\mathbf{F}^{T}\mathbf{F}|^{2}+|\mathbf{A}_{\varepsilon}^{T}(\theta)\mathbf{A}_{\varepsilon}(\theta)|^{2}\big{)}\leq 2c_{1}\varepsilon^{-1}\big{(}|\mathbf{F}|^{4}+|\mathbf{A}_{\varepsilon}(\theta)|^{4}).$ (87)
Let $M:=\sup_{\theta\in(\theta^{-},\theta^{+})}|\mathbf{A}(\theta)|^{4}$ and note that $\sup_{\theta\in\mathbb{R}}|\mathbf{A}_{\varepsilon}(\theta)|^{4}\leq M$ by the definition of $\mathbf{A}_{\varepsilon}(\theta)$ in (79). Moreover, $M$ is finite by the definition of $\mathbf{A}(\theta)$ in (8), since $|\mathbf{u}(\theta)|\leq|\mathbf{t}_{1}(\theta)|+|\mathbf{t}_{3}(\theta)|=|\mathbf{t}_{1}^{r}|+|\mathbf{t}_{3}^{r}|$ and $|\mathbf{v}(\theta)|\leq|\mathbf{t}_{2}(\theta)|+|\mathbf{t}_{4}(\theta)|=|\mathbf{t}_{2}^{r}|+|\mathbf{t}_{4}^{r}|$. Thus, the upper bound asserted is true for $C_{1,\varepsilon}=2c_{1}\varepsilon^{-1}\max\{1,M\}$.
For the lowerbound, set $\mathbf{f}_{i}:=\mathbf{F}\mathbf{e}_{i}$ and $\mathbf{a}_{i,{\varepsilon}}(\theta):=\mathbf{A}_{\varepsilon}(\theta)\mathbf{e}_{i}$ for $i=1,2$. We have
$W^{\varepsilon}_{1}(\theta,\mathbf{F})$ $=\frac{c_{1}}{j_{\varepsilon}(\mathbf{F})}\Big{|}\begin{pmatrix}|\mathbf{f}_{1}|^{2}&\mathbf{f}_{1}\cdot\mathbf{f}_{2}\\ \mathbf{f}_{1}\cdot\mathbf{f}_{2}&|\mathbf{f}_{2}|^{2}\end{pmatrix}-\begin{pmatrix}|\mathbf{a}_{1,\varepsilon}(\theta)|^{2}&\mathbf{a}_{1,\varepsilon}(\theta)\cdot\mathbf{a}_{2,\varepsilon}(\theta)\\ \mathbf{a}_{1,\varepsilon}(\theta)\cdot\mathbf{a}_{2,\varepsilon}(\theta)&|\mathbf{a}_{2,\varepsilon}(\theta)|^{2}\end{pmatrix}\Big{|}^{2}\geq\frac{c_{1}}{j_{\varepsilon}(\mathbf{F})}\sum_{i=1,2}\big{|}|\mathbf{f}_{i}|^{2}-|\mathbf{a}_{i,\varepsilon}(\theta)|^{2}\big{|}^{2}.$ (88)
Using Young’s inequality, we have for any $\delta>0$ that
$\sum_{i=1,2}\big{|}|\mathbf{f}_{i}|^{2}-|\mathbf{a}_{i,\varepsilon}(\theta)|^{2}\big{|}^{2}$ $=\sum_{i=1,2}\left(|\mathbf{f}_{i}|^{4}-2|\mathbf{f}_{i}|^{2}|\mathbf{a}_{i,\varepsilon}(\theta)|^{2}+|\mathbf{a}_{i,\varepsilon}(\theta)|^{4}\right)\geq\sum_{i=1,2}\left((1-\delta)|\mathbf{f}_{i}|^{4}+\big{(}1-\frac{1}{\delta}\big{)}|\mathbf{a}_{i,\varepsilon}(\theta)|^{4}\right).$ (89)
Taking $\delta=\frac{1}{2}$ and employing the definition of $M$ gives
$\sum_{i=1,2}\big{|}|\mathbf{f}_{i}|^{2}-|\mathbf{a}_{i,\varepsilon}(\theta)|^{2}\big{|}^{2}\geq\sum_{i=1,2}\big{(}\frac{1}{2}|\mathbf{f}_{i}|^{4}-\frac{1}{2}|\mathbf{a}_{i,\varepsilon}(\theta)|^{4}\big{)}\geq\frac{1}{4}|\mathbf{F}|^{4}-\frac{1}{2}M.$ (90)
It follows that
$W_{1}^{\varepsilon}(\theta,\mathbf{F})\geq\frac{c_{1}}{j_{\varepsilon}(\mathbf{F})}\big{(}\frac{1}{4}|\mathbf{F}|^{4}-\frac{1}{2}M\Big{)}\geq\frac{c_{1}}{4}\frac{|\mathbf{F}|^{4}}{|\mathbf{F}\mathbf{e}_{1}||\mathbf{F}\mathbf{e}_{2}|}-\frac{c_{1}}{2}\varepsilon^{-1}M\geq\frac{c_{1}}{4}|\mathbf{F}|^{2}-\frac{1}{2}c_{1}\varepsilon^{-1}M.$ (91)
Therefore, the desired lower bound holds with $c_{1,\varepsilon}:=\min\{\frac{c_{1}}{4},\tfrac{2}{c_{1}\varepsilon^{-1}M}\}$. ∎
We now develop similar results for $W_{2}^{\varepsilon}$ in (LABEL:eq:W12Eps).
Lemma 2.
There is a constant $C_{2,\varepsilon}>0$ such that
$0\leq W_{2}^{\varepsilon}(\theta,\mathbf{F},\mathcal{F})\leq C_{2,\varepsilon}|\mathcal{F}|^{2}\quad\text{ for all }(\theta,\mathbf{F},\mathcal{F})\in\mathbb{R}\times\mathbb{R}^{3\times 2}\times\mathbb{R}^{3\times 2\times 2}.$ (92)
Proof.
We only need to prove the upperbound. By standard estimates,
$W_{2}^{\varepsilon}(\theta,\mathbf{F},\mathcal{F})\leq 2c_{2}L_{\Omega}^{2}(\big{[}\mathbf{v}_{\varepsilon}(\theta)\cdot[\mathbf{v}^{\prime}]_{\varepsilon}(\theta)\big{]}^{2}+\big{[}\mathbf{u}_{\varepsilon}(\theta)\cdot[\mathbf{u}^{\prime}]_{\varepsilon}(\theta)\big{]}^{2})|\mathbf{K}_{\varepsilon}(\mathbf{F},\mathcal{F})|^{2}.$ (93)
Let $M_{\varepsilon}:=\sup_{\theta\in\mathbb{R}}(\big{[}\mathbf{v}_{\varepsilon}(\theta)\cdot[\mathbf{v}^{\prime}]_{\varepsilon}(\theta)\big{]}^{2}+\big{[}\mathbf{u}_{\varepsilon}(\theta)\cdot[\mathbf{u}^{\prime}]_{\varepsilon}(\theta)\big{]}^{2})$. This quantity is clearly finite by the smoothness of $\mathbf{u}(\theta),\mathbf{v}(\theta)$ on $(\theta^{-},\theta^{+})$ and by the definitions of their extensions $\mathbf{u}_{\varepsilon}(\theta),\mathbf{v}_{\varepsilon}(\theta),[\mathbf{u}^{\prime}]_{\varepsilon}(\theta)$ and $[\mathbf{v}^{\prime}]_{\varepsilon}(\theta)$ in (78) and (80). Standard estimates on $\mathbf{K}_{\varepsilon}(\mathbf{F},\mathcal{F})$ in (85) then yield
$|\mathbf{K}_{\varepsilon}(\mathbf{F},\mathcal{F})|^{2}\leq(|\mathcal{F}_{11}|^{2}+2|\mathcal{F}_{12}|^{2}+|\mathcal{F}_{22}|^{2})|\bm{\nu}_{\varepsilon}(\mathbf{F})|^{2}\leq|\mathcal{F}|^{2}$ (94)
where the last inequality uses that $|\mathcal{F}_{11}|^{2}+2|\mathcal{F}_{12}|^{2}+|\mathcal{F}_{22}|^{2}=|\mathcal{F}|^{2}$ and that $|\bm{\nu}_{\varepsilon}(\mathbf{F})|=1$ when $|\mathbf{F}\mathbf{e}_{1}\times\mathbf{F}\mathbf{e}_{2}|\geq\varepsilon$ and $<1$ otherwise. We conclude that $W_{2}^{\varepsilon}(\theta,\mathbf{F},\mathcal{F})\leq 2c_{2}L_{\Omega}^{2}M_{\varepsilon}|\mathcal{F}|^{2}$ and thus $C_{2,\varepsilon}:=2c_{2}L_{\Omega}^{2}M_{\varepsilon}$. ∎
We state one final result highlighting the key properties of $W^{\varepsilon}$ that go into proving Theorem 1.
Proposition 1.
There are positive constants $c_{\varepsilon},C_{\varepsilon}>0$ such that
$c_{\varepsilon}\Big{(}|\theta|^{2}+|\mathbf{g}|^{2}+|\mathbf{F}|^{2}+|\mathcal{F}|^{2}\Big{)}-\frac{1}{c_{\varepsilon}}\leq W^{\varepsilon}(\theta,\mathbf{g},\mathbf{F},\mathcal{F})\leq C_{\varepsilon}(|\theta|^{2}+|\mathbf{g}|^{2}+|\mathbf{F}|^{4}+|\mathcal{F}|^{2}+1)$ (95)
for all $(\theta,\mathbf{g},\mathbf{F},\mathcal{F})\in\mathbb{R}\times\mathbb{R}^{2}\times\mathbb{R}^{3\times 2}\times\mathbb{R}^{3\times 2\times 2}$. In addition, $W^{\varepsilon}(\theta,\mathbf{g},\mathbf{F},\mathcal{F})$ is continuous in all its arguments and convex in $(\mathbf{g},\mathcal{F})$ for each $(\theta,\mathbf{F})$.
Proof.
For the lowerbound, we have by (83) and Lemma’s 1-2 that
$W^{\varepsilon}(\theta,\mathbf{g},\mathbf{F},\mathcal{F})\geq c_{1,\varepsilon}|\mathbf{F}|^{2}-\frac{1}{c_{1,\varepsilon}}+d_{1}|\mathcal{F}|^{2}+d_{2}|\mathbf{g}|^{2}+d_{3}|\theta|^{2}.$ (96)
The choice $c_{\varepsilon}:=\min\{c_{1,\varepsilon},d_{1}L_{\Omega}^{2},d_{2}L_{\Omega}^{2},d_{3}\}$ gives the desired result. For the upperbound, we again have by (83) and Lemma’s 1-2 that
$W^{\varepsilon}(\theta,\mathbf{g},\mathbf{F},\mathcal{F})$ $\leq C_{1,\varepsilon}(|\mathbf{F}|^{4}+1)+d_{2}L_{\Omega}^{2}|\mathbf{g}|^{2}+(d_{1}L_{\Omega}^{2}+C_{2,\varepsilon})|\mathcal{F}|^{2}+d_{3}|\theta|^{2}$ (97)
$\leq(C_{1,\varepsilon}+d_{1}L_{\Omega}^{2}+d_{2}L_{\Omega}^{2}+d_{3}+C_{2,\varepsilon})(|\theta|^{2}+|\mathbf{g}|^{2}+|\mathbf{F}|^{4}+|\mathcal{F}|^{2}+1).$
Thus, $C_{\varepsilon}:=(C_{1,\varepsilon}+d_{1}L_{\Omega}^{2}+d_{2}L_{\Omega}^{2}+d_{3}+C_{2,\varepsilon})$ gives the desired upperbound.
As for the convexity statement, it is clear that $d_{1}L_{\Omega}^{2}|\mathcal{F}|^{2}+d_{2}L_{\Omega}^{2}|\mathbf{g}|^{2}$ is convex and quadratic in $(\mathbf{g},\mathcal{F})$, so the only questions concern $W_{2}^{\varepsilon}(\theta,\mathbf{F},\mathcal{F})$ in (LABEL:eq:W12Eps). However, $W_{2}^{\varepsilon}$ satisfies $W_{2}^{\varepsilon}(\theta,\mathbf{F},\mathcal{F})=c_{2}L_{\Omega}^{2}(\bm{\nu}_{\varepsilon}(\mathbf{F})\cdot\mathcal{F}\colon\mathbf{C}_{\varepsilon}(\theta))^{2}$ for all $(\theta,\mathbf{F},\mathcal{F})\in\mathbb{R}\times\mathbb{R}^{2\times 2}\times\mathbb{R}^{3\times 2\times 2}$, where $\mathbf{C}_{\varepsilon}(\theta):=\big{(}\mathbf{v}_{\varepsilon}(\theta)\cdot[\mathbf{v}^{\prime}]_{\varepsilon}(\theta)\big{)}\tilde{\mathbf{u}}_{0}\otimes\tilde{\mathbf{u}}_{0}+\big{(}\mathbf{u}_{\varepsilon}(\theta)\cdot[\mathbf{u}^{\prime}]_{\varepsilon}(\theta)\big{)}\tilde{\mathbf{v}}_{0}\otimes\tilde{\mathbf{v}}_{0}$. This function is clearly convex in $\mathcal{F}$. This completes the proof. ∎
Step 3. Energy estimates. We now go from pointwise estimates on the various energy densities that make up $E_{\text{int}}(\mathbf{y},\theta)$ to upper and lower bounds on the overall energy $E(\mathbf{y},\theta)$. For the coercivity result, we follow a presentation by Ball ([5], Theorem 7.6) for mixed boundary conditions.
Proposition 2.
Assume that $\Gamma_{\emph{n}}\subset\Gamma_{\emph{d}}$ are measurable, $\Gamma_{\emph{d}}$ has nonzero measure, and that $V^{\varepsilon}_{\Gamma}$ is non-empty. Let $\overline{\mathbf{b}}\in L^{2}(\Omega,\mathbb{R}^{3})$, $\overline{\mathbf{t}}\in L^{2}(\Gamma_{\emph{t}},\mathbb{R}^{3})$, $\overline{\mathbf{m}}\in L^{2}(\Gamma_{\emph{m}},\mathbb{R}^{3})$. Then, for all $(\mathbf{y},\theta)\in V^{\varepsilon}_{\Gamma}$, the following two statments hold: (1) $E(\mathbf{y},\theta)<+\infty$ and (2) there is a constant $c_{\varepsilon}^{\star}>0$ depending only on $c_{\varepsilon}$ from (95) and $\Omega$ such that
$E(\mathbf{y},\theta)\geq c_{\varepsilon}^{\star}\big{(}\|\theta\|^{2}_{H^{1}(\Omega)}+\|\mathbf{y}\|^{2}_{H^{2}(\Omega)}\big{)}-\frac{1}{c_{\varepsilon}^{\star}}\Big{(}\|\overline{\mathbf{b}}\|^{2}_{L^{2}(\Omega)}+\|\overline{\mathbf{t}}\|^{2}_{L^{2}(\Gamma_{\text{t}})}+\|\overline{\mathbf{m}}\|^{2}_{L^{2}(\Gamma_{\text{m}})}+\|\overline{\mathbf{y}}\|^{2}_{L^{2}(\Gamma_{\text{d}})}+1\Big{)}.$ (98)
Remark 1.
Since $V^{\varepsilon}_{\Gamma}$ is assumed to be non-empty above, $\|\overline{\mathbf{y}}\|_{L^{2}(\Gamma_{\emph{d}})}<\infty$ by the trace theorem for Sobolev maps. Thus, from the assumptions of Theorem 1 (the same as those in Proposition 2) and the inequality in (98), the infimum $E^{\star}:=\inf_{(\mathbf{y},\theta)\in V^{\varepsilon}_{\Gamma}}E(\mathbf{y},\theta)$ is not $\infty$- ∞.
Proof.
Fix any $(\mathbf{y},\theta)\in V^{\varepsilon}_{\Gamma}$ (which is non-empty by assumption) and note that $\theta\in H^{1}(\Omega)$ and $\mathbf{y}\in H^{2}(\Omega)$.
We first prove that $E(\mathbf{y},\theta)<+\infty$. The Cauchy-Schwarz inequality and the trace theorem provide an upper bound on $E_{\text{ext}}(\mathbf{y})$ in (27) of the form
$E_{\text{ext}}(\mathbf{y})$ $\leq\|\mathbf{y}\|_{L^{2}(\Omega)}\|\overline{\mathbf{b}}\|_{L^{2}(\Omega)}+\|T\mathbf{y}\|_{L^{2}(\Gamma_{\text{t}})}\|\overline{\mathbf{t}}\|_{L^{2}(\Gamma_{\text{t}})}+\|T\nabla\mathbf{y}\|_{L^{2}(\Gamma_{\text{m}})}\|\overline{\mathbf{m}}\|_{L^{2}(\Gamma_{\text{m}})}$ (99)
$\leq C_{0}\|\mathbf{y}\|_{H^{2}(\Omega)}\left(\|\overline{\mathbf{t}}\|_{L^{2}(\Gamma_{\text{t}})}+\|\overline{\mathbf{m}}\|_{L^{2}(\Gamma_{\text{m}})}+\|\overline{\mathbf{b}}\|_{L^{2}(\Omega)}\right)$
for some $C_{0}>0$ that only depends on $\Omega$, where $T\colon H^{1}(\Omega)\rightarrow L^{2}(\partial\Omega)$ in the first inequality is the trace operator mapping an $H^{1}$ function on $\Omega$ to its “trace”, an $L^{2}$ function on the boundary $\partial\Omega$ (see, for instance, [24]). We conclude that $E_{\text{ext}}(\mathbf{y})<\infty$ from the assumptions on the boundary fields $\overline{\mathbf{b}}$, $\overline{\mathbf{t}}$, and $\overline{\mathbf{m}}$ in the proposition and since $\mathbf{y}\in H^{2}(\Omega)$. Next, observe that, since $V^{\varepsilon}_{\Gamma}\subset V^{\varepsilon}$,
$E_{\text{int}}(\mathbf{y},\theta)=\int_{\Omega}W^{\varepsilon}(\theta,\nabla\theta,\nabla\mathbf{y},\nabla\nabla\mathbf{y})\mathop{}\!\mathrm{d}x\leq C_{\varepsilon}\Big{(}\|\theta\|_{H^{1}(\Omega)}^{2}+\|\nabla\mathbf{y}\|_{L^{4}(\Omega)}^{4}+\|\nabla\nabla\mathbf{y}\|_{L^{2}(\Omega)}^{2}+|\Omega|\Big{)}$ (100)
by the definition in (77) and the upperbound in (95) of Proposition 1. Since $\Omega\subset\mathbb{R}^{2}$ and $\mathbf{y}\in H^{2}(\Omega)$, the Sobolev embedding theorem implies that $\nabla\mathbf{y}\in L^{4}(\Omega)$. So it follows from (100) that $E_{\text{int}}(\mathbf{y},\theta)<+\infty$. Thus, $E(\mathbf{y},\theta):=E_{\text{int}}(\mathbf{y},\theta)+E_{\text{ext}}(\mathbf{y},\theta)<+\infty$, as desired.
We now prove the lowerbound in (98). Since $V^{\varepsilon}_{\Gamma}\subset V^{\varepsilon}$, we have by the definition of $W^{\varepsilon}$ in (77) and the lowerbound in (95) of Proposition 1 that
$E(\mathbf{y},\theta)\geq c_{\varepsilon}\int_{\Omega}\Big{(}|\theta|^{2}+|\nabla\theta|^{2}+|\nabla\mathbf{y}|^{2}+|\nabla\nabla\mathbf{y}|^{2}\Big{)}\mathop{}\!\mathrm{d}x-c_{\varepsilon}^{-1}|\Omega|+E_{\text{ext}}(\mathbf{y}).$ (101)
Focusing on $E_{\text{ext}}(\mathbf{y})$, applying a Cauchy–Schwarz inequality, a Young’s inequality for $\delta>0$, and a trace inequality leads to
$E_{\text{ext}}(\mathbf{y},\theta)$ $\geq-\delta\|\mathbf{y}\|_{L^{2}(\Omega)}\delta^{-1}\|\overline{\mathbf{b}}\|_{L^{2}(\Omega)}-\delta\|T\mathbf{y}\|_{L^{2}(\Gamma_{\text{t}})}\delta^{-1}\|\overline{\mathbf{t}}\|_{L^{2}(\Gamma_{\text{t}})}-\delta\|T\nabla\mathbf{y}\|_{L^{2}(\Gamma_{\text{m}})}\delta^{-1}\|\overline{\mathbf{m}}\|_{L^{2}(\Gamma_{\text{m}})}$ (102)
$\geq-\frac{\delta^{2}}{2}\Big{(}\|\mathbf{y}\|_{L^{2}(\Omega)}^{2}+\|T\mathbf{y}\|_{L^{2}(\Gamma_{\text{t}})}^{2}+\|T\nabla\mathbf{y}\|^{2}_{L^{2}(\Gamma_{\text{t}})}\Big{)}-\frac{1}{2\delta^{2}}\Big{(}\|\overline{\mathbf{b}}\|^{2}_{L^{2}(\Omega)}+\|\overline{\mathbf{t}}\|^{2}_{L^{2}(\Gamma_{\text{t}})}+\|\overline{\mathbf{m}}\|^{2}_{L^{2}(\Gamma_{\text{m}})}\Big{)}$
$\geq-\delta^{2}C_{1}\|\mathbf{y}\|_{H^{2}(\Omega)}^{2}-\frac{1}{\delta^{2}}\Big{(}\|\overline{\mathbf{b}}\|^{2}_{L^{2}(\Omega)}+\|\overline{\mathbf{t}}\|^{2}_{L^{2}(\Gamma_{\text{t}})}+\|\overline{\mathbf{m}}\|^{2}_{L^{2}(\Gamma_{\text{m}})}\Big{)}$
for some constant $C_{1}>0$ that depends only on $\Omega$. Furthermore, since $\Gamma_{\text{d}}$ has positive measure, a Poincaré type estimate (from [42], page 82) gives
[…]
for some constant $C_{2}>0$ that depends only on $\Omega$, where $\|\overline{\mathbf{y}}\|_{L^{2}(\Gamma_{\text{d}})}^{2}<\infty$ since $\overline{\mathbf{y}}\in H^{1/2}(\Gamma_{\text{d}},\mathbb{R}^{3})$. Combining the inequalities in (101-103) gives
$E(\mathbf{y},\theta)$ $\geq c_{\varepsilon}\int_{\Omega}\Big{(}|\mathbf{\theta}|^{2}+|\nabla\theta|^{2}+\lambda|\nabla\mathbf{y}|^{2}+|\nabla\nabla\mathbf{y}|^{2}\Big{)}\mathop{}\!\mathrm{d}x+c_{\varepsilon}(1-\lambda)C_{2}^{-1}\int_{\Omega}|\mathbf{y}|^{2}\mathop{}\!\mathrm{d}x-c_{\varepsilon}^{-1}|\Omega|$ (104)
$\qquad-\delta^{2}C_{1}\|\mathbf{y}\|_{H^{2}(\Omega)}^{2}-\frac{1}{\delta^{2}}\Big{(}\|\overline{\mathbf{b}}\|^{2}_{L^{2}(\Omega)}+\|\overline{\mathbf{t}}\|^{2}_{L^{2}(\Gamma_{\text{t}})}+\|\overline{\mathbf{m}}\|^{2}_{L^{2}(\Gamma_{\text{m}})}\Big{)}-c_{\varepsilon}(1-\lambda)\int_{\Gamma_{\text{d}}}|\overline{\mathbf{y}}|^{2}\mathop{}\!\mathrm{d}s$
for any choice of $\lambda\in(0,1)$ and $\delta>0$. By choosing such $\lambda$ and $\delta$ appropriately, we deduce from (104) that there is a constant $c_{\varepsilon}^{\star}>0$ that depends only on $c_{\varepsilon},C_{1},C_{2},$ and $\Omega$ (so $c_{\varepsilon}$ and $\Omega$) such that (98) holds. ∎
Step 4. Proof of Theorem 1. As the assumptions of the theorem are the same as that of Proposition 2, this proposition and Remark 1 establish that $E^{\star}:=\inf_{(\mathbf{y},\theta)\in V_{\Gamma}^{\varepsilon}}E(\mathbf{y},\theta)$ is finite. Hence, there exists a minimizing sequence $\{(\mathbf{y}_{j},\theta_{j})\}\subset V^{\varepsilon}_{\Gamma}$ such that $E(\mathbf{y}_{j},\theta_{j})\leq E^{\star}+\frac{1}{j}$ for all positive integers $j$. Moreover, the lowerbound in (98) of Proposition 2 furnishes the estimates on this sequence
$\|\theta_{j}\|^{2}_{H^{1}(\Omega)}+\|\mathbf{y}_{j}\|^{2}_{H^{2}(\Omega)}\leq\frac{1}{c_{\varepsilon}^{\star}}(E^{\star}+1)+\frac{1}{(c_{\varepsilon}^{\star})^{2}}\Big{(}\|\overline{\mathbf{b}}\|^{2}_{L^{2}(\Omega)}+\|\overline{\mathbf{t}}\|^{2}_{L^{2}(\Gamma_{\text{t}})}+\|\overline{\mathbf{m}}\|^{2}_{L^{2}(\Gamma_{\text{m}})}+\|\overline{\mathbf{y}}\|^{2}_{L^{2}(\Gamma_{\text{d}})}+1\Big{)}$ (105)
for a $c_{\varepsilon}^{\star}>0$ independent of $j$. As the bound on the right is finite (by the assumptions of the theorem) and independent of $j$, the sequence $\{(\mathbf{y}_{j},\theta_{j})\}$ is uniformly bounded in the $H^{2}\times H^{1}$ norm. It follows that there is a subsequence $\{(\mathbf{y}_{j_{k}},\theta_{j_{k}})\}$ that converges weakly in $H^{2}\times H^{1}$, i.e.,
$\mathbf{y}_{j_{k}}\rightharpoonup\mathbf{y}\text{ in }H^{2}(\Omega,\mathbb{R}^{3})\quad\text{ and }\quad\theta_{j_{k}}\rightharpoonup\theta\text{ in }H^{1}(\Omega,\mathbb{R}).$ (106)
We now show that $(\mathbf{y},\theta)$ belongs to $V_{\Gamma}^{\varepsilon}$, first by verifying that these fields satisfy the $\varepsilon$-inequalities contained in this set and then by verifying the boundary conditions. By Rellich’s theorem, there is subsequence (not relabeled) that satisfies $\nabla\mathbf{y}_{j_{k}}\rightarrow\nabla\mathbf{y}$ in $L^{2}(\Omega)$ and $\theta_{j_{k}}\rightarrow\theta$ in $L^{2}(\Omega)$. By this strong $L^{2}$-convergence, there is an even further subsequence (not relabeled) such that $\nabla\mathbf{y}_{j_{k}}(\mathbf{x})\rightarrow\nabla\mathbf{y}(\mathbf{x})$ and $\theta_{j_{k}}(\mathbf{x})\rightarrow\theta(\mathbf{x})$ for a.e. $\mathbf{x}\in\Omega$. Since $\theta_{j_{k}}(\mathbf{x})\geq\theta^{-}+\varepsilon$ for a.e. $\mathbf{x}\in\Omega$ for all $j_{k}$,
$\theta(\mathbf{x})=\lim_{k\rightarrow\infty}\theta_{j_{k}}(\mathbf{x})\geq\theta^{-}+\varepsilon$ (107)
for a.e. $\mathbf{x}\in\Omega$. Arguing similarly, we conclude that $\theta(\mathbf{x})\leq\theta^{+}-\varepsilon$ for a.e. $\mathbf{x}\in\Omega$. Finally, since the function $f(\mathbf{F}):=|\mathbf{F}\mathbf{e}_{1}\times\mathbf{F}\mathbf{e}_{2}|$ is continuous and since $f(\nabla\mathbf{y}_{j_{k}}(\mathbf{x}))=|\partial_{1}\mathbf{y}_{j_{k}}(\mathbf{x})\times\partial_{2}\mathbf{y}_{j_{k}}(\mathbf{x})|\geq\varepsilon$ for a.e. $\mathbf{x}\in\Omega$ for all $j_{k}$, we conclude that
$|\partial_{1}\mathbf{y}(\mathbf{x})\times\partial_{2}\mathbf{y}(\mathbf{x})|=f\Big{(}\lim_{k\rightarrow\infty}\nabla\mathbf{y}_{j_{k}}(\mathbf{x})\Big{)}=\lim_{k\rightarrow\infty}f(\nabla\mathbf{y}_{j_{k}}(\mathbf{x}))\geq\varepsilon$ (108)
for a.e. $\mathbf{x}\in\Omega$. We now verify the boundary conditions. Since the trace operator $T\colon H^{1}(\Omega)\rightarrow L^{2}(\partial\Omega)$ is bounded and linear, it is continuous and it is follows from (106) that
$T\mathbf{y}_{j_{k}}\rightharpoonup T\mathbf{y}\text{ in $L^{2}(\partial\Omega)$}\quad\text{ and }\quad T\nabla\mathbf{y}_{j_{k}}\rightharpoonup T\nabla\mathbf{y}\text{ in $L^{2}(\partial\Omega)$}.$ (109)
Since $T\mathbf{y}_{j_{k}}=\overline{\mathbf{y}}$ a.e. on $\Gamma_{\text{d}}$ and $T\nabla\mathbf{y}_{j_{k}}\mathbf{n}=\overline{\mathbf{s}}$ a.e. on $\Gamma_{\text{n}}$, we conclude from (109) that $T\mathbf{y}=\overline{\mathbf{y}}$ a.e. on $\Gamma_{\text{d}}$ and $T\nabla\mathbf{y}\mathbf{n}=\bar{\mathbf{s}}$ a.e. on $\Gamma_{\text{n}}$. Hence, $(\mathbf{y},\theta)\in V_{\Gamma}^{\varepsilon}$.
We are now ready to complete the proof. The properties established by Proposition 1 place us in a setting where we can directly apply a general result on weakly lower-semicontinuous functionals (Dacorogna [13], Theorem 3.23). In particular, we conclude from (106) that
$\liminf_{k\rightarrow\infty}E_{\text{int}}(\mathbf{y}_{j_{k}},\theta_{j_{k}})$ $=\liminf_{k\rightarrow\infty}\int_{\Omega}W^{\varepsilon}(\theta_{j_{k}},\nabla\theta_{j_{k}},\nabla\mathbf{y}_{j_{k}},\nabla\nabla\mathbf{y}_{j_{k}})\mathop{}\!\mathrm{d}x$ (110)
$\geq\int_{\Omega}W^{\varepsilon}(\theta,\nabla\theta,\nabla\mathbf{y},\nabla\nabla\mathbf{y})\mathop{}\!\mathrm{d}x=E_{\text{int}}(\mathbf{y},\theta),$
where the last equality follows since $(\mathbf{y},\theta)\in V_{\Gamma}^{\varepsilon}$. The basic reasoning behind this inequality is that $W^{\varepsilon}$ is continuous in the lower order terms $(\theta,\nabla\mathbf{y})$ and convex in these terms gradients $(\nabla\theta,\nabla\nabla\mathbf{y})$. Next, we note that the potential $E_{\text{ext}}(\mathbf{y}_{j_{k}})$ associated to applied forces converges trivially to $E_{\text{ext}}(\mathbf{y})$ given (106) and (109) since the terms $\mathbf{y}_{j_{k}}$, $T\mathbf{y}_{j_{k}}$ and $T\nabla\mathbf{y}_{j_{k}}$ are paired with functions $\overline{\mathbf{b}}$, $\overline{\mathbf{t}}$ and $\overline{\mathbf{m}}$ in their dual space, i.e.,
$\lim_{k\rightarrow\infty}E_{\text{ext}}(\mathbf{y}_{j_{k}})=E_{\text{ext}}(\mathbf{y}).$ (111)
[…]
$E^{\star}\geq\liminf_{k\rightarrow\infty}E(\mathbf{y}_{j_{k}},\theta_{j_{k}})=\liminf_{k\rightarrow\infty}\big{(}E_{\text{int}}(\mathbf{y}_{j_{k}},\theta_{j_{k}})+E_{\text{ext}}(\mathbf{y}_{j_{k}})\big{)}\geq E_{\text{int}}(\mathbf{y},\theta)+E_{\text{ext}}(\mathbf{y})=E(\mathbf{y},\theta).$ (112)
Since $E^{\star}=\inf_{V_{\Gamma}^{\varepsilon}}E$ and $(\mathbf{y},\theta)\in V_{\Gamma}^{\varepsilon}$, the inequalities above are actually equalities, i.e., $E^{\star}=E(\mathbf{y},\theta)$, and thus $(\mathbf{y},\theta)$ is a minimizer. This completes the proof of the existence of minimizers.
Appendix B On the derivation of the governing equations
Here, we derive a key identity used to obtain the strong form of the equilibrium equations and natural boundary conditions from weak form, namely, that
$\int_{\Omega}\Big{\{}\mathbf{P}(\mathbf{y},\theta)\colon\nabla\mathbf{w}+\big{\langle}\mathcal{H}(\mathbf{y},\theta),\nabla\nabla\mathbf{w}\big{\rangle}+q(\mathbf{y},\theta)\eta+\mathbf{j}(\theta)\cdot\nabla\eta\Big{\}}\mathop{}\!\mathrm{d}x$ (113)
$\quad=\int_{\Omega}\big{(}\operatorname{div}\big{[}\operatorname{div}\mathcal{H}(\mathbf{y},\theta)-\mathbf{P}(\mathbf{y},\theta)\big{]}\big{)}\cdot\mathbf{w}\mathop{}\!\mathrm{d}x+\int_{\partial\Omega}\big{[}\mathbf{P}(\mathbf{y},\theta)-\nabla\mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{I}+\mathbf{n}^{\perp}\otimes\mathbf{n}^{\perp})\big{]}\mathbf{n}\cdot\mathbf{w}\mathop{}\!\mathrm{d}s$
$\qquad\quad+\int_{\partial\Omega}\big{(}\mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{n}\otimes\mathbf{n})\big{)}\cdot(\nabla\mathbf{w})\mathbf{n}\mathop{}\!\mathrm{d}s+\int_{\Omega}\big{(}q(\mathbf{y},\theta)-\nabla\cdot\mathbf{j}(\theta)\big{)}\eta\mathop{}\!\mathrm{d}x+\int_{\partial\Omega}\big{(}\mathbf{j}(\theta)\cdot\mathbf{n}\big{)}\eta\mathop{}\!\mathrm{d}s.$
holds for all sufficiently smooth $\mathbf{y},\mathbf{w}\colon\Omega\rightarrow\mathbb{R}^{3}$ and $\theta,\eta\colon\Omega\rightarrow(\theta^{-},\theta^{+})$, where $\mathbf{n}$ is the outward normal to $\partial\Omega$ and $\mathbf{n}^{\perp}$ is the corresponding unit tangent vector. (See also below (39) for any questions about the tensor notation.)
To begin, use integration by parts and the divergence theorem to the term on the left in (LABEL:eq:importForStrongForm) to obtain
$\int_{\Omega}\Big{\{}\mathbf{P}(\mathbf{y},\theta)\colon\nabla\mathbf{w}+\big{\langle}\mathcal{H}(\mathbf{y},\theta),\nabla\nabla\mathbf{w}\big{\rangle}+q(\mathbf{y},\theta)\eta+\mathbf{j}(\theta)\cdot\nabla\eta\Big{\}}\mathop{}\!\mathrm{d}x$ (114)
$\qquad=\int_{\Omega}\big{(}\operatorname{div}\big{[}\operatorname{div}\mathcal{H}(\mathbf{y},\theta)-\mathbf{P}(\mathbf{y},\theta)\big{]}\big{)}\cdot\mathbf{w}\mathop{}\!\mathrm{d}x+\int_{\partial\Omega}\big{[}\mathbf{P}(\mathbf{y},\theta)-\operatorname{div}\mathcal{H}(\mathbf{y},\theta)\big{]}\mathbf{n}\cdot\mathbf{w}\mathop{}\!\mathrm{d}s$
$\qquad\quad+\int_{\partial\Omega}\big{(}\mathcal{H}(\mathbf{y},\theta)\cdot\mathbf{n}\big{)}\colon\nabla\mathbf{w}\mathop{}\!\mathrm{d}s+\int_{\Omega}\big{(}q(\mathbf{y},\theta)-\nabla\cdot\mathbf{j}(\theta)\big{)}\eta\mathop{}\!\mathrm{d}x+\int_{\partial\Omega}\big{(}\mathbf{j}(\theta)\cdot\mathbf{n}\big{)}\eta\mathop{}\!\mathrm{d}s$
where $[\mathcal{H}\cdot\mathbf{n}]_{i\alpha}=[\mathcal{H}]_{i\alpha\beta}n_{\beta}$. Next, write $\nabla\mathbf{w}=\partial_{\mathbf{n}}\mathbf{w}\otimes\mathbf{n}+\partial_{\mathbf{n}^{\perp}}\mathbf{w}\otimes\mathbf{n}^{\perp}$ and observe that
$\int_{\partial\Omega}\big{(}\mathcal{H}(\mathbf{y},\theta)\cdot\mathbf{n}\big{)}\colon\nabla\mathbf{w}\mathop{}\!\mathrm{d}s=\int_{\partial\Omega}\Big{\{}\big{(}\mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{n}\otimes\mathbf{n})\big{)}\cdot\partial_{\mathbf{n}}\mathbf{w}+\big{(}\mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{n}\otimes\mathbf{n}^{\perp})\big{)}\cdot\partial_{\mathbf{n}^{\perp}}\mathbf{w}\Big{\}}\mathop{}\!\mathrm{d}s.$ (115)
Another round of integration by parts on the last term in (115) furnishes
$\int_{\partial\Omega}\big{(}\mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{n}\otimes\mathbf{n}^{\perp})\big{)}\cdot\partial_{\mathbf{n}^{\perp}}\mathbf{w}\mathop{}\!\mathrm{d}s=-\int_{\partial\Omega}\partial_{\mathbf{n}^{\perp}}\big{(}\mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{n}\otimes\mathbf{n}^{\perp})\big{)}\cdot\mathbf{w}\mathop{}\!\mathrm{d}s.$ (116)
In particular, $\int_{\partial\Omega}\partial_{\mathbf{n}^{\perp}}\big{[}\big{(}\mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{n}\otimes\mathbf{n}^{\perp})\big{)}\cdot\mathbf{w}\big{]}\mathop{}\!\mathrm{d}s=0$ because the tangential derivative of a field integrated on a closed boundary always vanishes. To finish, note that $\partial_{\mathbf{n}^{\perp}}\mathbf{n}=\kappa\mathbf{n}^{\perp}$, where $\kappa$ is the curvature of $\partial\Omega$. Likewise, $\partial_{\mathbf{n}^{\perp}}\mathbf{n}^{\perp}=-\kappa\mathbf{n}$. Thus,
$\partial_{\mathbf{n}^{\perp}}\big{[}\mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{n}\otimes\mathbf{n}^{\perp})\big{]}$ $=\partial_{\mathbf{n}^{\perp}}\big{[}\mathcal{H}(\mathbf{y},\theta)\big{]}\colon(\mathbf{n}\otimes\mathbf{n}^{\perp})+\kappa\mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{n}^{\perp}\otimes\mathbf{n}^{\perp}-\mathbf{n}\otimes\mathbf{n})$ (117)
$=\big{[}\nabla\mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{n}^{\perp}\otimes\mathbf{n}^{\perp})\big{]}\mathbf{n},$
where the term proportional to $\kappa$ in the first line vanishes since $\mathcal{H}$ is symmetric ($[\mathcal{H}]_{i\alpha\beta}=[\mathcal{H}]_{i\beta\alpha}$) and $(\mathbf{n}^{\perp}\otimes\mathbf{n}^{\perp}-\mathbf{n}\otimes\mathbf{n})$ is skew symmetric. We complete the derivation by substituting the results of (115-117) into (LABEL:eq:importForStrongForm1) and using that $\operatorname{div}\mathcal{H}=\nabla\mathcal{H}\colon\mathbf{I}$ and $\partial_{\mathbf{n}}\mathbf{w}=(\nabla\mathbf{w})\mathbf{n}$.
Appendix C Derivation of the consistency term
This appendix derives the consistency terms introduced in the interior penalty method in Section 4.2. The basic idea is that, if our numerical method happens upon an exact solution to the governing equations in (40), then it should identify these fields as solutions to the discrete numerical problem as well. This demand leads naturally to the form of $\mathscr{A}_{\text{con}}^{h}$ in (47). We refer to Section 4.1 for any questions about the notation in this derivation.
Let $(\mathbf{y}_{h},\theta_{h})\in V_{\Gamma_{\text{d}}}^{h}$ and $(\mathbf{w}_{h},\eta_{h})\in V_{0}^{h}$. Observe using integration by parts that $\mathscr{A}_{0}^{h}$ in (45) satisfies
$\mathscr{A}_{0}^{h}((\mathbf{y}_{h},\theta_{h});(\mathbf{w}_{h},\eta_{h}))$ $=\sum_{T\in\mathcal{T}^{h}}\bigg{\{}\int_{T}\operatorname{div}\big{[}\operatorname{div}\mathcal{H}(\mathbf{y}_{h},\theta_{h})-\mathbf{P}(\mathbf{y}_{h},\theta_{h})\big{]}\cdot\mathbf{w}_{h}\mathop{}\!\mathrm{d}x$ (118)
$\qquad+\int_{\partial T}\big{[}\big{(}\mathbf{P}(\mathbf{y}_{h},\theta_{h})-\operatorname{div}\mathcal{H}(\mathbf{y}_{h},\theta_{h})\big{)}\mathbf{n}\cdot\mathbf{w}_{h}+\big{(}\mathcal{H}(\mathbf{y}_{h},\theta_{h})\cdot\mathbf{n}\big{)}\colon\nabla\mathbf{w}_{h}\big{]}\mathop{}\!\mathrm{d}s$
$\qquad+\int_{T}\big{(}q(\mathbf{y}_{h},\theta_{h})-\nabla\cdot\mathbf{j}(\theta_{h})\big{)}\eta_{h}\mathop{}\!\mathrm{d}x+\int_{\partial T}\big{(}\mathbf{j}(\theta_{h})\cdot\mathbf{n}\big{)}\eta_{h}\mathop{}\!\mathrm{d}s\bigg{\}}.$
This integral can be reorganized by replacing the boundary integrals over each $\partial T$ of the triangulation $\mathcal{T}^{h}$ with an identical set of integrals along the edges $e\in\mathcal{E}^{h}$, leading to
$\mathscr{A}_{0}^{h}((\mathbf{y}_{h},\theta_{h});(\mathbf{w}_{h},\eta_{h}))$ $=\sum_{T\in\mathcal{T}^{h}}\int_{T}\Big{\{}\operatorname{div}\big{[}\operatorname{div}\mathcal{H}(\mathbf{y}_{h},\theta_{h})-\mathbf{P}(\mathbf{y}_{h},\theta_{h})\big{]}\cdot\mathbf{w}_{h}+\big{(}q(\mathbf{y}_{h},\theta_{h})-\nabla\cdot\mathbf{j}(\theta_{h})\big{)}\eta_{h}\Big{\}}\mathop{}\!\mathrm{d}x$ (119)
$\qquad+\sum_{e\in\mathcal{E}_{\text{int}}^{h}}\int_{e}\Big{\{}\llbracket\mathbf{P}(\mathbf{y}_{h},\theta_{h})-\operatorname{div}\mathcal{H}(\mathbf{y}_{h},\theta_{h})\rrbracket\mathbf{n}_{e}\cdot\mathbf{w}_{h}+\big{(}\{\mskip-5.0mu\{\mathcal{H}(\mathbf{y}_{h},\theta_{h})\}\mskip-5.0mu\}\cdot\mathbf{n}_{e}\big{)}\colon\llbracket\nabla\mathbf{w}_{h}\rrbracket\Big{\}}\mathop{}\!\mathrm{d}s$
$\qquad+\int_{\Gamma_{\text{t}}}\Big{\{}\big{(}\mathbf{P}(\mathbf{y}_{h},\theta_{h})-\operatorname{div}\mathcal{H}(\mathbf{y}_{h},\theta_{h})\big{)}\mathbf{n}\cdot\mathbf{w}_{h}\Big{\}}\mathop{}\!\mathrm{d}s+\int_{\partial\Omega}\big{(}\mathcal{H}(\mathbf{y}_{h},\theta_{h})\cdot\mathbf{n}\big{)}\colon\nabla\mathbf{w}_{h}\mathop{}\!\mathrm{d}s$
$\qquad+\sum_{e\in\mathcal{E}_{\text{int}}^{h}}\int_{e}\big{(}\llbracket\mathbf{j}(\theta_{h})\rrbracket\cdot\mathbf{n}_{e}\big{)}\eta_{h}\mathop{}\!\mathrm{d}s+\int_{\partial\Omega}\big{(}\mathbf{j}(\theta_{h})\cdot\mathbf{n}\big{)}\eta_{h}\mathop{}\!\mathrm{d}s.$
The jump conditions on the interior edges of the mesh are due to the lack of $H^{2}$-regularity in the deformation $\mathbf{y}_{h}$ and test field $\mathbf{w}_{h}$ in our discrete formulation. However, an exact solution $(\mathbf{y},\theta)\in V_{\Gamma}$ to the strong form of the governing equation in (40) satisfies
$\llbracket\mathbf{P}(\mathbf{y},\theta)-\operatorname{div}\mathcal{H}(\mathbf{y},\theta)\rrbracket\mathbf{n}_{e}=\mathbf{0}\quad\text{ and }\quad\llbracket\mathbf{j}(\theta)\rrbracket\cdot\mathbf{n}_{e}=0$ (120)
on each $e\in\mathcal{E}_{\text{int}}^{h}$. Thus, replacing $(\mathbf{y}_{h},\theta_{h})$ in (119) with an exact solution and subtracting off the boundary terms $\mathscr{B}_{0}(\mathbf{w}_{h})$ gives
$\mathscr{A}_{0}^{h}((\mathbf{y},\theta);(\mathbf{w}_{h},\eta_{h}))-\mathscr{B}_{0}(\mathbf{w}_{h})=\sum_{e\in\mathcal{E}_{\text{int}}^{h}}\int_{e}\big{(}\{\mskip-5.0mu\{\mathcal{H}(\mathbf{y},\theta)\}\mskip-5.0mu\}\cdot\mathbf{n}_{e}\big{)}\colon\llbracket\nabla\mathbf{w}_{h}\rrbracket\mathop{}\!\mathrm{d}s+\int_{\Gamma_{\text{n}}}\big{(}\mathcal{H}(\mathbf{y},\theta)\cdot\mathbf{n}\big{)}\colon\nabla\mathbf{w}_{h}\mathop{}\!\mathrm{d}s.$ (121)
The right-hand side here is actually $-\mathscr{A}_{\text{con}}^{h}((\mathbf{y},\theta);(\mathbf{w}_{h},\eta_{h}))$ for $\mathscr{A}_{\text{con}}^{h}$ defined in (47). In adding this term to left side of our discrete formulation (along with $\mathscr{A}_{\text{sta}}^{h}$ and $\mathscr{A}_{\text{bnd}}^{h}$), it follows that
$\mathscr{A}^{h}((\mathbf{y},\theta);(\mathbf{w}_{h},\eta_{h}))=\mathscr{B}_{0}(\mathbf{w}_{h})\quad\text{ for all }(\mathbf{w}_{h},\eta_{h})\in V^{h}_{0}$ (122)
for any $(\mathbf{y},\theta)\in V_{\Gamma}$ that solves (40), since $\mathscr{A}_{\text{sta}}^{h}((\mathbf{y},\theta);(\mathbf{w}_{h},\eta_{h}))$ and $\mathscr{A}_{\text{bnd}}^{h}((\mathbf{y},\theta),(\mathbf{w}_{h},\eta_{h}))$ vanish trivially on an exact solution. This is the desired consistency.
Appendix D Initialization of the boundary conditions
For investigating large deformations and angle fields in our numerical framework, the Newton solver is improved when we initialize it with a pair of fields $(\mathbf{y}^{0}_{h},\theta^{0}_{h})$ that a priori satisfy the boundary conditions. We therefore perform two preliminary (and much simpler) numerical calculations to initialized such fields.
We start by solving the minimization problem $\min\{\int_{\Omega}\frac{1}{2}|\nabla\nabla\mathbf{y}|^{2}\mathop{}\!\mathrm{d}x+E_{\text{ext}}(\mathbf{y})\colon\mathbf{y}=\overline{\mathbf{y}}\text{ on }\Gamma_{\text{d}}\;(\nabla\mathbf{y})\mathbf{n}=\overline{\mathbf{s}}\text{ on}\Gamma_{\text{n}}\}$ using $C^{0}$ finite elements and the interior penalty method. Referencing again the notation in Section 4.1 and mimicking the exposition in Section 4.2 and Appendix C, this problem is formulated in the weak form as: Find a $\mathbf{y}^{0}_{h}\in\widetilde{V}_{\Gamma_{\text{d}}}^{h}$ such that
$\mathscr{A}_{\text{pre}}^{h}(\mathbf{y}^{0}_{h};\mathbf{w}_{h})=\mathscr{B}_{\text{pre}}^{h}(\mathbf{w}_{h})\quad\text{ for all }\mathbf{w}_{h}\in\widetilde{V}^{h}_{0},$ (123)
where $\widetilde{V}^{h}_{\Gamma_{\text{d}}}$ and $\widetilde{V}^{h}_{0}$ are the approximation spaces given by restricting $V_{\Gamma_{\text{d}}}^{h}$ and $V^{h}_{0}$, respectively, to the deformation variable (i.e., neglecting the angle variable). In this setting, $\mathscr{A}_{\text{pre}}^{h}$ is a bilinear form defined by
$\mathscr{A}_{\text{pre}}^{h}(\mathbf{y}^{0}_{h};\mathbf{w}_{h})$ $:=\sum_{T\in\mathcal{T}^{h}}\int_{T}\langle\nabla_{h}\nabla\mathbf{y}_{h}^{0},\nabla_{h}\nabla\mathbf{w}_{h}\rangle\mathop{}\!\mathrm{d}x-\sum_{e\in\mathcal{E}_{\text{int}}^{h}}\int_{e}\big{(}\{\mskip-5.0mu\{\nabla_{h}\nabla\mathbf{y}^{0}_{h}\}\mskip-5.0mu\}\cdot\mathbf{n}_{e}\big{)}\colon\llbracket\nabla\mathbf{w}_{h}\rrbracket\mathop{}\!\mathrm{d}s$ (124)
$\qquad-\sum_{e\in\mathcal{E}_{\text{int}}^{h}}\int_{e}\big{(}\{\mskip-5.0mu\{\nabla_{h}\nabla\mathbf{w}_{h}\}\mskip-5.0mu\}\cdot\mathbf{n}_{e}\big{)}\colon\llbracket\nabla\mathbf{y}_{h}^{0}\rrbracket\mathop{}\!\mathrm{d}s+\sum_{e\in\mathcal{E}_{\text{int}}^{h}}\frac{\alpha}{|e|}\int_{e}\llbracket\nabla\mathbf{y}^{0}_{h}\rrbracket:\llbracket\nabla\mathbf{w}_{h}\rrbracket\mathop{}\!\mathrm{d}s$
$\qquad-\int_{\Gamma_{\text{n}}}\big{(}\nabla_{h}\nabla\mathbf{y}^{0}_{h}\colon(\mathbf{n}\otimes\mathbf{n})\big{)}\cdot\nabla\mathbf{w}_{h}\mathbf{n}\mathop{}\!\mathrm{d}s-\int_{\Gamma_{\text{n}}}\big{(}\nabla_{h}\nabla\mathbf{w}_{h}\colon(\mathbf{n}\otimes\mathbf{n})\big{)}\cdot(\nabla\mathbf{y}^{0}_{h})\mathbf{n}\mathop{}\!\mathrm{d}s$
$\qquad+\sum_{e\in\mathcal{E}^{h}_{\Gamma_{\text{n}}}}\frac{\alpha}{|e|}\int_{e}(\nabla\mathbf{y}_{h}^{0})\mathbf{n}\cdot(\nabla\mathbf{w}_{h})\mathbf{n}\mathop{}\!\mathrm{d}s$
The linear form $\mathscr{B}_{\text{pre}}^{h}$ on the right side of (123) is defined by
$\mathscr{B}^{h}_{\text{pre}}(\mathbf{w}_{h}):=\mathscr{B}_{0}(\mathbf{w}_{h})-\int_{\Gamma_{\text{n}}}\big{(}\nabla_{h}\nabla\mathbf{w}_{h}\colon(\mathbf{n}\otimes\mathbf{n})\big{)}\cdot\bar{\mathbf{s}}\mathop{}\!\mathrm{d}s+\sum_{e\in\mathcal{E}^{h}_{\Gamma_{\text{n}}}}\frac{\alpha}{|e|}\int_{e}\overline{\mathbf{s}}\cdot(\nabla\mathbf{w}_{h})\mathbf{n}\mathop{}\!\mathrm{d}s.$ (125)
In both $\mathscr{A}_{\text{pre}}^{h}$ and $\mathscr{B}^{h}_{\text{pre}}$, we include symmetry terms that mirror $(\mathbf{y}^{0}_{h};\mathbf{w}_{h})\equiv(\mathbf{w}_{h};\mathbf{y}^{0}_{h})$ in the argument for $\mathscr{A}_{\text{pre}}^{h}$. These terms are standard (see [16, 20]) and easy to implement when using the interior penalty method to solve a linear PDE because the weak form of the internal energy is a symmetric bilinear form. Their inclusion, while not necessary, can improve the effectiveness and efficiency of the numerical solver. In the typical initialization step to obtain a $\mathbf{y}_{h}^{0}$, the value of the penalty parameter is chosen as $\alpha=20$ in an effort to strongly impose the slope boundary conditions $\nabla\mathbf{y}_{h}^{0}\mathbf{n}=\overline{\mathbf{s}}$; it is subsequently reduced to $\alpha=0.1$ when solving (44).
Given a $\mathbf{y}_{h}^{0}$ solving the weak form above, we produce an initial guess $\theta^{0}_{h}$ for the actuation field by approximating the metric constraint. Specifically we compute $\theta^{0}_{h}$ as
$\theta^{0}_{h}:=\operatorname*{argmin}_{\theta_{h}\in W^{h}}\Big{\{}\int_{\Omega}\left\lvert(\nabla\mathbf{y}^{0}_{h})^{T}\nabla\mathbf{y}^{0}_{h}-\mathbf{A}^{T}(\theta_{h})\mathbf{A}(\theta_{h})\right\rvert^{2}\mathop{}\!\mathrm{d}x+d_{2}\int_{\Omega}\left\lvert\nabla\theta_{h}\right\rvert^{2}\mathop{}\!\mathrm{d}x\Big{\}}$ (126)
in the space $W^{h}:=\{\theta_{h}\in C^{0}(\Omega,\mathbb{R})\colon\theta_{h}|_{T}\in P^{1}(T)\;\;\forall\;\;T\in\mathcal{T}^{h}\}$, where the second term with $d_{2}>0$ is the regularizing term in (22) associated to the gradient of actuation. As (126) has a standard finite element implementation, we do not detail it here for the sake of brevity.
Appendix E Computation of consistent reaction forces
Here we explain how to compute reaction forces in the simulations. To start, consider a model problem where $(\mathbf{y},\theta)$ solve the governing equations in (40) so that the weak form $\mathscr{A}_{0}((\mathbf{y},\theta);(\mathbf{w},\eta))=\mathscr{B}_{0}(\mathbf{w})$ holds for all $(\mathbf{w},\theta)\in V_{0}$ (see the beginning of Section 3.4 for the definition of $V_{0}$ and Section 4.2 for the definitions of $\mathscr{A}_{0}$ and $\mathscr{B}_{0}$). The reaction forces emerge when we replace the test functions $\mathbf{w}$ in the weak form with a generic test function $\mathbf{w}_{\text{f}}\in H^{2}(\Omega,\mathbb{R}^{3})$ that does not satisfy the vanishing Dirichlet boundary conditions contained in $V_{0}$. Indeed, using the identity in (LABEL:eq:importForStrongForm) and that $(\mathbf{y},\theta)$ solve (40), this replacement yields
$\mathscr{A}_{0}((\mathbf{y},\theta);(\mathbf{w}_{\text{f}},0))$ $=\int_{\Omega}\overline{\mathbf{b}}\cdot\mathbf{w}_{\text{f}}\mathop{}\!\mathrm{d}x+\int_{\Gamma_{\text{t}}}\overline{\mathbf{t}}\cdot\mathbf{w}_{\text{f}}\mathop{}\!\mathrm{d}s+\int_{\Gamma_{\text{d}}}\big{[}\big{(}\mathbf{P}(\mathbf{y},\theta)-\nabla\mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{I}+\mathbf{n}^{\perp}\otimes\mathbf{n}^{\perp})\big{)}\mathbf{n}\big{]}\cdot\mathbf{w}_{\text{f}}\mathop{}\!\mathrm{d}s$ (127)
$\qquad+\int_{\Gamma_{\text{m}}}\overline{\mathbf{m}}\cdot(\nabla\mathbf{w}_{\text{f}})\mathbf{n}\mathop{}\!\mathrm{d}s+\int_{\Gamma_{\text{n}}}\big{(}\mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{n}\otimes\mathbf{n})\big{)}\cdot(\nabla\mathbf{w}_{\text{f}})\mathbf{n}\mathop{}\!\mathrm{d}s$
The definition of $\mathscr{B}_{0}(\mathbf{w})$ then allows us to conclude that
$\mathscr{A}_{0}((\mathbf{y},\theta);(\mathbf{w}_{\text{f}},0))-\mathscr{B}_{0}(\mathbf{w}_{\text{f}})$ $=\int_{\Gamma_{\text{d}}}\big{[}\big{(}\mathbf{P}(\mathbf{y},\theta)-\nabla\mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{I}+\mathbf{n}^{\perp}\otimes\mathbf{n}^{\perp})\big{)}\mathbf{n}\big{]}\cdot\mathbf{w}_{\text{f}}\mathop{}\!\mathrm{d}s$ (128)
$\qquad+\int_{\Gamma_{\text{n}}}\big{(}\mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{n}\otimes\mathbf{n})\big{)}\cdot(\nabla\mathbf{w}_{\text{f}})\mathbf{n}\mathop{}\!\mathrm{d}s.$
Let us now specialize to the pinching boundary conditions of Section 5.3, where this method is employed. The Dirichlet boundary set is $\Gamma_{\text{d}}=\Gamma_{L}\cup\Gamma_{R}$ and $\Gamma_{\text{n}}=\emptyset$. Clearly, the force on $\Gamma_{L}$ in the $\mathbf{e}_{1}$-direction is balanced by that on $\Gamma_{R}$. So we compute the boundary force in the plots by choosing a test function such that $\mathbf{w}_{\text{f}}=\mathbf{0}$ on $\Gamma_{L}$ and $\mathbf{w}_{\text{f}}=\mathbf{e}_{1}$ on $\Gamma_{R}$. It follows that
$\mathscr{A}_{0}((\mathbf{y},\theta);(\mathbf{w}_{\text{f}},0))-\mathscr{B}_{0}(\mathbf{w}_{\text{f}})=\int_{\Gamma_{\text{R}}}\big{[}\big{(}\mathbf{P}(\mathbf{y},\theta)-\nabla\mathcal{H}(\mathbf{y},\theta)\colon(\mathbf{I}+\mathbf{n}^{\perp}\otimes\mathbf{n}^{\perp})\big{)}\mathbf{n}\big{]}\cdot\mathbf{e}_{1}\mathop{}\!\mathrm{d}s,$ (129)
which is the desired total generalized force on the right boundary.
[…]